### v2

# references_lipschitz_polytopes.bib

Abstract.
In 2010, Vershik proposed a new combinatorial invariant of metric spaces given by a class of polytopes that arise in the theory of optimal transport and are called “Wasserstein polytopes” or “Kantorovich-Rubinstein polytopes” in the literature. Answering a question posed by Vershik, we describe the stratification of the metric cone induced by the combinatorial type of these polytopes through a hyperplane arrangement. Moreover, we study its relationships with the stratification by combinatorial type of the injective hull (i.e., the tight span) and, in particular, with certain types of metrics arising in phylogenetic analysis. We also compute enumerative invariants in the case of metrics on up to six points.
In memory of Andreas W. M. Dress.
1. Introduction
This article is centered around the study of finite metric spaces through the lens of polytopes and hyperplane arrangements. Finite metric spaces arise in several applied contexts. Let us mention for instance mathematical biology, where finite metrics model genetic dissimilarities between different species [Phylogenetics]. In this setting, a main research direction is to identify suitable classes of metric spaces and study the combinatorics and geometry of the associated subset of the metric cone, e.g., for geometric statistics. This field of research goes back to the study of phylogenetic trees [Bandelt-Dress-Trees, BHV], has grown to include more general phylogenetic networks [HRS] and is presently very active (see e.g., [Devadoss-petti, Francis_Steel, Fischer_Francis]). A combinatorial invariant of metric spaces that is widely used in the aforementioned applications as well as for theoretical considerations is their injective hull, introduced by Isbell [Isbell] and rediscovered by Dress [Dress] under the name tight span. The stratification of the metric cone by combinatorial type of tight spans is equivalent to the one determined by the subdivisions of the second hypersimplex studied by Sturmfels and Yu [SturmfelsYu] and recently by Casabella, Joswig, and Kastner [CJK24].
Motivated by the theory of optimal transport, Vershik described a correspondence between finite metric spaces and a class of symmetric convex polytopes, the so-called Kantorovich-Rubinstein-Wasserstein (KRW) polytopes or fundamental polytopes [Vershik].
Definition 1.1.
The KRW polytope of an $n$-metric $\rho$ is a polytope in $\mathbb{R}^{n}$ defined as the convex hull:
$KRW(\rho)=\operatorname{conv}\left\{\left.\frac{e_{i}-e_{j}}{\rho_{ij}}\right|1\leq i,j\leq n\right\},$
where $e_{i}$ is the $i$-th standard basis vector of $\mathbb{R}^{n}$. From now on, we abbreviate
$p_{ij}:=\frac{e_{i}-e_{j}}{\rho_{ij}}\textrm{ for }i\neq j.$
The metric is called generic if it is strict and the KRW polytope is simplicial.
If $\rho$ is the $n$-metric with $\rho_{ij}=1$ for all $i\neq j$, then $KRW(\rho)$ is the type $A_{n}$ root polytope. If $\rho$ is a graph-metric (i.e., there is a graph $G$ with vertex set $[n]$ such that $\rho_{ij}$ is the number of edges of a shortest path from $i$ to $j$ in $G$), then $KRW(\rho)$ is the symmetric edge polytope of $G$ [dali-delucchi-michalek], also known as the adjacency polytope of $G$ and, as such, appearing in different applied contexts [Chen, CDM]. If $\rho$ is a tree metric (i.e., it can be represented as the distances among a subset of the vertices of a metric tree), the number of faces of $KRW(\rho)$ can be computed explicitly, and the dual polytope is a zonotope [DelucchiHoessly].
The dual of a KRW polytope is the so-called Lipschitz polytope [Vershik]. This is a symmetric alcoved polytope [LP1]; a class of polytopes that recently gained significant attention due to its connections to theoretical physics in the form of positive geometries and polypositroids [LP2].
The problem of understanding the combinatorial structure (e.g., computing face numbers) of KRW and Lipschitz polytopes is open and significant for applications, see e.g., [Becedasetal, venturelloetal]. In this article, we focus on a related problem that goes back to a question by Vershik.
Problem ([Vershik, ”General Problem”, §1]).
Study and classify finite metric spaces according to the combinatorial properties of their KRW polytopes.
The first progress on this question was achieved by Gordon and Petrov who gave a description of the face poset of KRW polytopes via linear inequalities on the values of the metric [GordonPetrov], see Section 2.2 for further details. This is the starting point for an in-depth study of KRW polytopes using the theory of hyperplane arrangements initiated in this paper.
1.1. The main results
We now summarize the main contributions of this article:
  1. (1)
    We introduce a Wasserstein arrangement whose cells correspond to the combinatorial types of KRW polytopes. This arrangement builds upon the Gordon–Petrov description of the faces of KRW polytopes and could be of independent interest. Moreover, via the above connection to alcoved polytopes, intersecting this arrangement with the metric cone naturally yields the type-fan of symmetric alcoved polytopes.
  2. (2)
    Using the computer algebra system OSCAR [OSCAR-book] together with the Julia package CountingChambers.jl [BEK23] and the software TOPCOM [Rambau] we obtain an enumeration of the combinatorial types of KRW polytopes of generic metrics on $n=4,5,6$ points. This is displayed in Table 1, where the number of combinatorial types of generic KRW polytopes – i.e., the number of chambers of the Wasserstein arrangement – is shown in the column titled “Labeled”. The column “Unlabeled” shows the number of combinatorial types up to combinatorial isomorphisms (where we are allowed to permute the vertices to obtain a combinatorial isomorphism), which is the number of chambers of the Wasserstein arrangement modulo the natural action of the symmetric group $S_{n}$ via coordinate permutation.
    This table already shows that the subdivision of the metric cone by combinatorial types of KRW polytopes is much finer than the subdivision by tight spans as for $n=6$ there are $194,160$ different tight spans which up to symmetry come in $339$ orbits [SturmfelsYu].
    $n$ Unlabeled Labeled
    3 1 1
    4 1 6
    5 12 882
    6 25,224 6,677,863,200
    Table 1. Combinatorial types of KRW polytopes of generic $n$-metrics.
  3. (3)
    We clarify the relation between KRW polytopes and tight spans (Definition 5.3) by providing examples of five-point metrics with isomorphic tight spans and combinatorially different KRW polytopes and vice versa. Hence, the two fan structures of the metric fan induced by the tight spans or the KRW polytopes are not refinements of each other.
  4. (4)
    We compare the fan structure induced by the Wasserstein arrangement on the metric cone with some types of metrics that have appeared prominently in the literature on phylogenetic analysis. In particular, we show that the space of tree-like metrics and of circular-decomposable metrics are subfans of the Wasserstein fan (Proposition 6.1, Theorem 6.4), thus so is the simplicial fan of Circular Split Networks introduced by Devadoss and Petti [Devadoss-petti] (and appearing also in work of Hacking, Keel and Tevelev [HKT]), see Remark 6.5. Moreover, the set of metrics satisfying the six-point condition of Dress-Huber-Koolen-Moulton [SixPoints] is a subfan of the Wasserstein fan as well (Proposition 6.16). However, the set of all totally split-decomposable metrics is not a subfan of the Wasserstein fan (see Example 6.13), and neither is the set of all consistent, totally split-decomposable metrics (Remark 6.17). We leave the following as open questions.
    Question 1.2 (Tree-based metrics).
    The notion of tree-based phylogenetic network has gathered considerable attention in the literature on phylogenetic analysis, since its introduction by Francis and Steel in 2015 [Francis_Steel, Francis_Semple_Steel]. The space of such tree-based phylogenetic networks has been the object of recent research by Fischer and Francis [Fischer_Francis]. In our context, it is therefore natural to ask whether the metric spaces arising from tree-based phylogenetic networks are a union of strata of the Wasserstein fan.
    Question 1.3 (Antipodal metrics).
    A further type of metrics that has been studied by Dress, Huber and Moulton is the class of antipodal metrics [DHMantipodal]. Antipodality imposes strong constraints on the combinatorial structure of the tight span (e.g., a finite metric space is antipodal if and only if its tight span has a unique maximal cell, see [HuKoMo1]). It is then natural to ask whether antipodal metric spaces form a subfan of the Wasserstein fan.
  5. (5)
    We derive formulae for the number of vertices and edges of the KRW polytopes of non-generic strict metrics. To the best of our knowledge, the following question is open.
    Question 1.4.
    Compute the full $f$-vector of KRW polytopes of non-generic strict metrics. Can the approach and the formulae given in Section 7 be generalized?
1.2. Roadmap of the article
We introduce the main objects used throughout in Section 2. The Wasserstein arrangement, capturing the combinatorial structures of KRW polytopes, is defined and discussed in Section 3. In Section 4 we explain the connection between KRW polytopes and a certain class of regular triangulations of type $A$ root polytopes. Subsequently, Section 5 is devoted to the relation between KRW polytopes and tight spans. We explore several noteworthy subfans of the Wasserstein arrangement stemming from phylogeny in Section 6. We present results concerning the $f$-vector of KRW polytopes associated with strict metrics in Section 7. In Section 8, we summarize our computational findings on KRW polytopes for metrics on up to $6$ points. We discuss the relations between the tight span and a certain subcomplex that we call the “pruned tight span” in Appendix A. Lastly, in Appendix B we summarize our computational findings by listing the types of KRW polytopes of strict metric spaces on $n=5$ points along with some enumerative invariants.
Acknowledgments
We thank Nick Early and Benjamin Schröter for fruitful discussions, as well as Laura Casabella and Michael Joswig for helpful comments on earlier versions of this manuscript. We are grateful to Jörg Rambau for providing us with relevant triangulations of the root polytope computed using TOPCOM.
LK and LM are supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – SFB-TRR 358/1 2023 – 491392403 and SPP 2458 – 539866293.
2. The main characters
2.1. The metric cone
We begin by defining finite metric spaces and setting up some terminology.
Definition 2.1.
An $n$-metric is a real symmetric $n\times n$ matrix $\rho$ with entries $\rho_{ij}$ for $1\leq i,j\leq n$ satisfying
  1. (1)
    $\rho_{ii}=0$ for all $1\leq i\leq n$,
  2. (2)
    $\rho_{ij}>0$ for all $1\leq i\neq j\leq n$ and
  3. (3)
    $\rho_{ij}+\rho_{jk}\geq\rho_{ik}$ for all $1\leq i,j,k\leq n$.
The metric $\rho$ is called strict if $\rho_{ij}+\rho_{jk}>\rho_{ik}$ for all $j\in\left[n\right]\setminus\{i,k\}$.
Every $n$-metric $\rho$ is given as a symmetric $n\times n$ matrix with zero diagonal, thus it is determined by the $\binom{n}{2}$ values $\rho_{ij}$ with $i<j$. This establishes a bijective correspondence between the set of all $n$-metrics and the following subset of $\mathbb{R}^{\binom{n}{2}}$.
Definition 2.2.
Consider the vector space $\mathbb{R}^{\binom{n}{2}}$ with coordinates $x_{\{i,j\}}$ indexed by pairs of elements of $[n]$. The metric cone on $n$ elements is the subset $\mathbf{M}_{n}\subseteq\mathbb{R}^{\binom{n}{2}}$ defined by
$x_{\{i,j\}}>0,\quad x_{\{i,j\}}+x_{\{j,k\}}\geq x_{\{i,k\}},\textrm{ for all pairwise distinct }i,j,k\in[n]$
. The pseudometric cone, denoted by $\overline{\mathbf{M}}_{n}$, is the (topological) closure of $\mathbf{M}_{n}$.
Remark 2.3.
The space $\overline{\mathbf{M}}_{n}$ is a polyhedral cone. It contains points that do not correspond to metrics but rather to pseudometrics (where $\rho_{ij}=0$ can be allowed). The set of all strict metrics on $n$ elements is the interior of $\mathbf{M}_{n}$.
Example 2.4 (Split pseudometrics).
Let $A\uplus B$ be a nontrivial bipartition of $[n]$. The pseudometric $\delta_{A,B}$ with
$(\delta_{A,B})_{i,j}:=\min\{|A\cap\{i,j\}|,|B\cap\{i,j\}|\}$
has value 00 when $i$ and $j$ are in the same part of the bipartition and has value $1$ otherwise. Any metric of this form is called an $n$-split, the partition itself $\sigma=A\uplus B$ is called a split.
2.2. Admissible graphs and strict metrics
Let $\rho$ be an $n$-metric and $F$ any face of the polytope $KRW(\rho)$. We associate with the face $F$ a directed graph $G(F)$ on the vertex set $[n]$ which contains the edge $(i,j)$ if the point $p_{ij}$ lies on $F$. Following [GordonPetrov], the collection of graphs of the form $G(F)$ is called the combinatorial structure of $KRW(\rho)$.
Theorem 2.5 ([GordonPetrov], Theorem 3).
Let $\rho$ be an $n$-metric and $G=([n],E)$ be a directed graph with the set of vertices $[n]$. The following are equivalent:
  1. (1)
    There exists a facet $F$ of the polytope $KRW(\rho)$ containing every vertex $p_{ij}$ for $(i,j)\in~{}E$.
  2. (2)
    For any array of directed edges $(x_{i},y_{i}),1\leq i\leq k$ of $G$ with all $x_{i}$ pairwise distinct and all $y_{i}$ pairwise distinct:
    (2.1) $\sum_{i=1}^{k}\rho_{x_{i}y_{i}}\leq\sum_{i=1}^{k}\rho_{x_{i}y_{i+1}},$
    where $y_{k+1}=y_{1}$.
Definition 2.6.
A directed graph $G$ on the vertex set $[n]$ is called admissible for an $n$-metric $\rho$ if it satisfies one of the two equivalent conditions of Theorem 2.5.
In general, the distinctness of the $x_{i}$ and $y_{i}$ does not mean that the two sets are disjoint. We have $x_{i}=y_{j}$ for some $i$ and $j$ if and only if the sequence $(x_{j},y_{j})$, $(x_{i},y_{i})$ is a directed path of length $2$ in $G$. However, for certain metrics it suffices to consider disjoint sets of start- and endvertices.
Lemma 2.7.
If $\rho$ be a strict metric, the condition in Theorem 2.5.(2) can only be satisfied if $x_{1},\ldots,x_{k}$, $y_{1},\ldots,y_{k}$ are pairwise distinct.
Proof.
Let $\rho$ be strict and consider an instance of Theorem 2.5.(2) with $x_{i}=y_{j}$ for some $i$ and $j$. Then (2.1) applied to $(x_{j},y_{j})$, $(x_{i},y_{i})=(y_{j},y_{i})$ gives
$\rho_{xy}+\rho_{yz}\leq\rho_{yy}+\rho_{xz}=\rho_{xz}$
and thus, $\rho_{xy}+\rho_{yz}=\rho_{xz}$ which contradicts the assumption on $\rho$ being strict. ∎
Corollary 2.8.
If $\rho$ is strict, admissible graphs do not contain any directed path with more than $2$ vertices.
We close this section with a characterization of generic metric spaces (in the sense of Definition 1.1).
Theorem 2.9 ([GordonPetrov], Theorem 5).
A strict $n$-metric $\rho$ is generic if and only if for arbitrary $2k$ distinct points $x_{1},\dots,x_{k},y_{1},\dots,y_{k}$ in $\left[n\right]$ the minimum of the terms
$\sum_{i=1}^{k}\rho_{x_{i},y_{\pi(i)}}$
is attained by a unique permutation $\pi\in S_{k}$.
2.3. Polyhedra and their combinatorial type
We refer to [ZieglerLoP] for basics and terminology about polyhedra and only recall some fundamental definitions. A polyhedron is any subset of Euclidean space that is obtained intersecting a finite number of closed halfspaces. A polytope is a bounded polyhedron – equivalently, the convex hull of a finite set of points. Every polyhedron is topologically closed. The affine hull of a polytope $P$ is the smallest affine subspace containing $P$, and the relative interior of $P$ is the interior of $P$ as a subset of its affine hull. If $P$ is any convex set in Euclidean space, we call face of $P$ any intersection of $P$ with the boundary of a closed halfspace containing $P$.
A polyhedral complex is a family $\mathscr{K}$ of polyhedra (all in the same ambient space) that contains the empty polyhedron and all faces of each of its members, and such the intersection of any two members of $\mathscr{K}$ is a face of both. The face poset of $\mathscr{K}$ is the set $\mathscr{K}$ partially ordered by inclusion.
Remark 2.10 (Labeled and unlabeled combinatorial type of polyhedral complexes).
We say that two finite polyhedral complexes $\mathscr{K},\mathscr{K^{\prime}}$ have the same combinatorial type (or ”are combinatorially isomorphic”, written $\mathscr{K}\simeq\mathscr{K^{\prime}}$) if there is an abstract polyhedral isomorphism (i.e., a face-preserving PL-homeomorphism) between them. If some or all of the cells of $\mathscr{K}$ and $\mathscr{K^{\prime}}$ are labeled using the same set of labels, and there is an abstract polyhedral isomorphism that restricts to a label-preserving bijection on the set of labeled faces, then we say that $\mathscr{K}$ and $\mathscr{K^{\prime}}$ have the same ”labeled” combinatorial type, and write $\mathscr{K}\simeq_{\ell}\mathscr{K^{\prime}}$.
Note that $\mathscr{K}\simeq\mathscr{K^{\prime}}$ if and only if the posets of faces of $\mathscr{K}$ and $\mathscr{K^{\prime}}$ are isomorphic [ZieglerLoP, §2.2], and $\mathscr{K}\simeq_{\ell}\mathscr{K^{\prime}}$ is equivalent to the existence of such an isomorphism of posets of faces that preserves the labels.
Example 2.11.
If $\rho$ is an $n$-metric, the polytope $KRW(\rho)$ together with all its faces defines a polyhedral complex, whose vertices are naturally labeled by ordered pairs of distinct elements of $[n]$ (i.e., if $p_{ij}$ is a vertex of $KRW(\rho)$, it is labeled by $(i,j)$). By Section 2.2, the KRW polytopes of two metrics $\rho_{1}$, $\rho_{2}$ have the same labeled combinatorial type if and only if the set of admissible labeled graphs for $\rho_{1}$ concides with the set of admissible labeled graphs for $\rho_{2}$, where the labeling of the graphs is induced by the labeling of the edges $(i,j)$, which are in one-to-one correspondence with the vertices $p_{ij}$.
We will call cone any subset $\mathbf{C}\subseteq R^{d}$ that contains every linear combination with positive coefficients of any of its finite subsets. Note that we thus slightly deviate from the standard terminology where zero coefficients are allowed, as in [ZieglerLoP, §1.1] in order to preserve the term ”metric cone” for $\mathbf{M}_{n}$. We call polyhedral cone any polyhedron containing 00 that is also a cone (this is equivalent to the standard notion).
Following standard terminology, we call polyhedral fan any polyhedral complex all of whose members are polyhedral cones, see [ZieglerLoP, §7.1]. We will call fan any set of (not necessarily polyhedral) cones that contains all faces of each of its members, and such the intersection of any two members is a face of both: again, this is not standard but allows us, e.g., to speak of the ”Wasserstein fan” $\mathscr{F}_{n}$ in Definition 3.3.
3. The Wasserstein arrangement and the Wasserstein fan
We now define an arrangement of hyperplanes whose faces encode the combinatorial structures of KRW polytopes.
Definition 3.1 (Wasserstein arrangement).
Let $n$ be a positive integer. We define a hyperplane arrangement in $\mathbb{R}^{\binom{n}{2}}$ as follows. Given $1<k$ and $k$-tuples $\mathbf{a}=(a_{1},\ldots,a_{k})$, $\mathbf{b}=(b_{1},\ldots,b_{k})$ with $a_{i},b_{i}\in[n]$ for all $i=1,\ldots,k$, define a hyperplane
$H_{\mathbf{a,b}}:=\left\{x\in\mathbb{R}^{\binom{n}{2}}\left|\sum_{i=1}^{k}x_{\{a_{i},b_{i}\}}=\sum_{i=1}^{k}x_{\{a_{i},b_{i+1}\}}\right.\right\}\quad\textrm{where }b_{k+1}=b_{1}.$
The “positive side” of $H_{\mathbf{a,b}}$ is
$H_{\mathbf{a,b}}^{+}:=\left\{x\in\mathbb{R}^{\binom{n}{2}}\left|\sum_{i=1,\ldots,k}x_{\{a_{i},b_{i}\}}\leq\sum_{i=1,\ldots,k}x_{\{a_{i},b_{i+1}\}}\right.\right\}.$
The Wasserstein arrangement is then the set of hyperplanes
$\mathcal{W}_{n}:=\left\{H_{\mathbf{a,b}}\left|\begin{array}[]{l}1<k\leq n,\,\,\mathbf{a,b}\in[n]^{k}\\ a_{1},\ldots,a_{k},b_{1},\ldots,b_{k}\textrm{ mutually distinct}\end{array}\right.\right\}.$
Remark 3.2.
In the case $k=2$, the tuple $(a_{1},a_{2})$ forms the same hyperplane when combined with either of the tuples $(b_{1},b_{2})$ or $(b_{2},b_{1})$. In order for the notion of positive side to be well-defined, we establish the convention that we chose the tuple starting with $\min\{b_{1},b_{2}\}$.
Definition 3.3 (Wasserstein fan).
Let $\mathscr{F}_{n}$ denote the fan determined by the intersections of the Wasserstein arrangement $\mathcal{W}_{n}$ with the metric cone $\mathbf{M}_{n}$. We call $\mathscr{F}_{n}$ the Wasserstein fan. Moreover, write ¯$\overline{\mathscr{F}}_{n}$ for the (polyhedral) fan of pseudometrics, defined by the intersections of $\mathcal{W}_{n}$ with all faces of the cone $\overline{\mathbf{M}}_{n}$.
Remark 3.4.
The metric cone $\mathbf{M}_{n}$ intersects every face of the Wasserstein arrangement $\mathcal{W}_{n}$. Indeed, let $p$ be any point on some face of $\mathcal{W}_{n}$. The point $p+k\mathbf{1}$ for some $k\geq 0$ and $\mathbf{1}$ the all-ones vector in $\mathbb{R}^{\binom{n}{2}}$ lies on exactly the same hyperplanes of the arrangement $\mathcal{W}_{n}$ as the point $p$ since the defining equations involve the same number of variables on both sides of the equation. But for $k$ large enough, we observe that the point $p+k\mathbf{1}$ is in $\mathbf{M}_{n}$.
Therefore, the arrangement $\mathcal{W}_{n}$ and the fan $\mathscr{F}_{n}$ have the same number of $i$-cells for any $i$. In particular, we can count the number of maximal cells of the fan $\mathscr{F}_{n}$ by counting the number of chambers of the arrangement $\mathcal{W}_{n}$.
Example 3.5.
The Wasserstein arrangement $\mathcal{W}_{4}$ in $\mathbb{R}^{\binom{4}{2}}$ consists of the three hyperplanes given by the equations
$H_{(1,2),(3,4)}:$ $\quad x_{\{1,3\}}+x_{\{2,4\}}=x_{\{1,4\}}+x_{\{2,3\}},$
$H_{(1,3),(2,4)}:$ $\quad x_{\{1,2\}}+x_{\{3,4\}}=x_{\{1,4\}}+x_{\{2,3\}},$
$H_{(1,4),(2,3)}:$ $\quad x_{\{1,2\}}+x_{\{3,4\}}=x_{\{1,3\}}+x_{\{2,4\}}.$
This arrangement has a $4$-dimensional lineality space, see Lemma 3.11. Figure 7 represents $\mathcal{W}_{4}$ modulo its lineality space.
Remark 3.6.
Each hyperplane $H_{\mathbf{a,b}}$ corresponds to a cycle $C_{H_{\mathbf{a,b}}}$ of the complete graph $K_{n}$ determined by the sequence of vertices $a_{1},b_{1},a_{k},b_{k},\ldots,a_{2},b_{2}$ (see Figure 1 for an example with $k=3$). The edges in the cycle have alternating orientations (all vertices have indegree either 0 or 2) and thus, the two sides of the hyperplane $H_{\mathbf{a,b}}$ determine which set of edges of the cycle is admissible (the even or the odd edges of the cycle). Similar to the halfspace notation, we refer to the set of edges $\{(a_{i},b_{i})_{1\leq i\leq k}\}$ by $C_{H_{\mathbf{a,b}}}^{+}$ and write $C_{H_{\mathbf{a,b}}}^{-}$ for the other set.
The indexing by the pair of $k$-tuples will be necessary in order to specify a well-defined positive side.
Figure 1. A hyperplane $H_{\mathbf{a,b}}$ with $k=3$ interpreted as cycle.
Proposition 3.7.
The arrangement $\mathcal{W}_{n}$ consists of
$|\mathcal{W}_{n}|=\frac{1}{2}\sum_{k=2}^{\lfloor\frac{n}{2}\rfloor}\binom{n}{2k}(2k-1)!$
many hyperplanes.
Proof.
As discussed in Remark 3.6, each hyperplane in $\mathcal{W}_{n}$ is determined by an oriented cycle in in the complete graph $K_{n}$ of even length at least four. Such cycles arise by choosing $2k$ vertices amongst the $n$ vertices and then choosing a cyclic permutation of length $2k$. There are $(2k-1)!$ such permutations. As reversing the orientation of the cycle yields the same hyperplane, we count every hyperplane exactly twice. ∎
Remark 3.8.
To study the combinatorial types of strict metrics, it suffices to consider the intersection of the interior of $\mathbf{M}_{n}$ with $\mathcal{W}_{n}$, since by Corollary 2.8 admissible graphs for strict metrics cannot contain directed paths with more than one edge.
Definition 3.9.
Recall Example 2.4. The pseudometrics of the form $\delta_{\{i\},[n]\setminus\{i\}}\in\mathbb{R}^{\binom{n}{2}}$ for $i=1,\ldots,n$ are called elementary splits.
Let $W$ denote the linear span of the elementary splits and let ¯$\overline{\mathscr{E}}_{n}:=\overline{\mathbf{M}}_{n}\cap W$ be the set of all pseudometrics that are linear combination of elementary splits. Accordingly, we write $\mathscr{E}_{n}:=\mathbf{M}_{n}\cap W$ for the cone of metrics that are combinations of elementary splits.
Remark 3.10.
The set of elementary splits is “compatible” in the sense of [BandeltDress] (e.g., because the metric obtained as the sum of all elementary splits is represented by a tree with $n$ leaves, each at distance $1$ from a unique inner vertex), and thus linearly independent by [BandeltDress, Corollary 4]. Therefore, $\dim(W)=n$ and ¯$\overline{\mathscr{E}}_{n}$ is an $n$-dimensional simplicial cone given as the convex hull of the rays generated by the elementary splits.
Lemma 3.11.
The lineality space of the Wasserstein arrangement $\mathcal{W}_{n}$ is the $n$-dimensional subspace $W$ spanned by the elementary splits (i.e., $W=\bigcap\mathcal{W}_{n}$).
Proof.
Let $L:=\bigcap\mathcal{W}_{n}$. We first prove that $W\subseteq L$. If a metric $\rho$ is in a hyperplane $H_{\mathbf{a,b}}$ and we define
$\rho^{\prime}:=\rho+\lambda\cdot\delta_{\{i\},[n]\setminus\{i\}}\in\mathbb{R}^{\binom{n}{2}}$
for $\lambda>0$ and some $i$, then the $\lambda$ appears in an equal number of summands on both sides of the defining equation of $H_{\mathbf{a,b}}$ and therefore, $\rho^{\prime}\in H_{\mathbf{a,b}}$. Thus, $W$ is contained in $H_{\mathbf{a,b}}$ for all $\mathbf{a}$ and $\mathbf{b}$ and is thus contained in the lineality space of the arrangement $\mathcal{W}_{n}$.
Now it is enough to prove that $\dim(W)\geq\dim(L)$. Since $\dim(W)=n$ (Remark 3.10), it is enough to find ${\binom{n}{2}}-n$ hyperplanes in $\mathcal{W}_{n}$ whose normal vectors are linearly independent.
For this, identify the elements of the metric space $[n]$ with the vertices of a regular $n$-gon and consider the set $D$ of all pairs of distinct elements of $[n]$ that define a diagonal (i.e., not an edge) of the $n$-gon. Thus $|D|=\binom{n}{2}-n$. For $\{i,j\}\in D$ consider $H_{(i,j+1),(j,i+1)}$ (we understand the sum modulo $n$, so that $n+1=1$), with normal vector $v_{ij}=e_{\{i,j\}}+e_{\{i+1,j+1\}}-e_{\{i,i+1\}}-e_{\{j,j+1\}}$.
We claim that the set $\{v_{ij}\mid\{i,j\}\in D\}$ is a linearly independent family. In order to do so, we arrange these vectors as columns of a matrix. Consider the following total ordering on the set of all pairs of distinct elements of $[n]$:
$\{1,2\}\prec\{2,3\}\prec\dots\prec\{n,1\}\prec\{1,3\}\prec\{2,4\}\prec\dots\prec\{n,2\}\prec\{1,3\}\prec\{2,4\}\prec\dots$
Under the bijection of unordered pairs with sides and diagonals of a regular $n$-gon, this corresponds to ordering these segments in increasing order of length and, for sides and diagonals of the same length $k$, in the cyclic order starting with the diagonal $\{1,k\}$.
Let $A$ be the $\binom{n}{2}\times|D|$-matrix whose columns are the vectors $v_{ij}$ listed in increasing $\prec$-order of $\{i,j\}$, and where the $k$-th row corresponds to the coordinate $e_{\{i,j\}}$ with $\{i,j\}$ in $k$-th place with respect to the ordering $\prec$. Then $A$ has the following block-form:
$A=\left(\begin{array}[]{c|c|c|c}W_{1}&W_{2}&W_{3}&\cdots\\ \hline\cr A_{1}&0&0&\cdots\\ 0&A_{2}&0&\cdots\\ 0&0&A_{3}&\cdots\\ \vdots&&&\ddots\\ \end{array}\right);\quad\textrm{ with }A_{j}=\begin{pmatrix}1&0&0&\cdots\\ 1&1&0&0&\cdots\\ 0&1&1&0&\cdots\\ 0&0&1&1&\ddots\\ \vdots&&\ddots&\ddots&\ddots\end{pmatrix},$
and the matrices $W_{i}$ have $n$ rows. Since clearly each $A_{i}$ has full rank, so does $A$. This proves that the $v_{ij}$ are linearly independent and thus that
$\dim(L)\leq\dim\left(\bigcap_{\{i,j\}\in D}H_{\{i,j\},\{i+1,j+1\}}\right)={\binom{n}{2}}-|D|=n.\qed$
The next proposition relates the combinatorial structure of KRW polytopes with the face structure of the Wasserstein fan.
Proposition 3.12.
Let $n\geq 2$ be an integer and consider two $n$-metrics $\rho^{(1)},\rho^{(2)}$. The following two statements are equivalent:
  1. (1)
    The metrics $\rho^{(1)},\rho^{(2)}$ are in the open part of one cone $\sigma$ of the Wasserstein fan $\mathscr{F}_{n}$.
  2. (2)
    The combinatorial structures of the polytopes $KRW(\rho^{(1)})$ and $KRW(\rho^{(2)})$ are the same, i.e., both polytopes have the same set of admissible directed graphs.
Proof.
Following Theorem 2.5, the sign pattern of a metric on an open cone in the fan $\mathscr{F}_{n}$, i.e., describing whether the metric lies on, “under” or “above” a hyperplane, specifies the collection of admissible graphs of the metric. As the latter is the combinatorial structure of the KRW polytope of the metric, the statement follows. ∎
Remark 3.13.
The hyperplanes $H_{\mathbf{a,b}}$ in $\mathcal{W}_{n}$ were already mentioned as ”exceptional planes” in the proof of Theorem 7 in [GordonPetrov]. The arrangement $\mathcal{A}\mathcal{B}_{n}$ of [Tran, Definition 26] is an arrangement of hyperplanes in the space $\mathbb{R}^{n^{2}-n}$ with coordinates $y_{(i,j)}$ indexed by ordered pairs of distinct elements of $[n]$. The arrangement $\mathcal{W}_{n}$ can be thought of as the restriction of $\mathcal{A}\mathcal{B}_{n}$ to the symmetric subspace defined by $y_{(i,j)}=y_{(j,i)}$ for all $i\neq j$.
4. Regular triangulations and the secondary fan
Following [DLRS] we start by recalling the definition of a regular subdivision of a set of points in $\mathbb{R}^{n}$, in particular of a convex polytope.
Definition 4.1.
Let $\mathcal{P}=\{a_{1},\dots,a_{k}\}$ be set of points in $\mathbb{R}^{n}$. A point $\omega\in\mathbb{R}^{k}$ defines a regular subdivision $\mathcal{P}(\omega)$ as follows: A subset $\{{i_{1}},\dots,{i_{r}}\}$ is a facet of $\mathcal{P}(\omega)$ if there exists a vector $c\in\mathbb{R}^{n}$ such that
(4.1) $\begin{cases}c\cdot a_{j}=\omega_{j}&\text{ if }j\in\{i_{1},\dots,i_{r}\}\text{ and}\\ c\cdot a_{j}<\omega_{j}&\text{ if }j\notin\{i_{1},\dots,i_{r}\}.\\ \end{cases}$
The subdivision $\mathcal{P}(\omega)$ is called a regular triangulation if all facets of the subdivision are simplices. This happens if the vector $\omega$ is sufficiently generic.
Geometrically, a regular subdivision arises by “lifting” the points in $\mathcal{P}$ to $\mathbb{R}^{n+1}$ where we use the coordinates of $\omega$ as heights. The faces of $\mathcal{P}(\omega)$ are then the “lower faces” of the convex hull of the lifted points, where a lower face is a face that has a normal vector with negative last coordinate.
Definition 4.2.
Let $\mathcal{R}_{n}^{0}=\{e_{i}-e_{j}\mid i,j\in\left[n\right]\}$. This is the set of vertices of the (full) root polytope of type $A$ together with the origin.
The next proposition establishes a relationship between regular subdivisions of $\mathcal{R}_{n}^{0}$ and KRW polytopes. A similar statement also appeared in the second arXiv version of the preprint [JS19].
Proposition 4.3.
Let $\rho$ be an $n$-metric.
  1. (1)
    The metric $\rho$ induces a regular subdivision $\mathcal{R}_{n}^{0}(\rho)$ of $\mathcal{R}_{n}^{0}$ by assigning the height $\rho_{ij}$ to the points $e_{i}-e_{j}$ and the height 00 to the origin.
  2. (2)
    The subdivision $\mathcal{R}_{n}^{0}(\rho)$ is point-symmetric with respect to the origin and every maximal cell contains the origin.
  3. (3)
    For any index set $I\subseteq[n]^{2}$ the following are equivalent:
    1. (i)
      The set $\{p_{ij}\mid(i,j)\in I\}$ is the set of all vertices contained in a facet of $KRW(\rho)$.
    2. (ii)
      The set $\{e_{i}-e_{j}\mid(i,j)\in I\}\cup\{0\}$ is a maximal cell of the regular subdivision $\mathcal{R}_{n}^{0}(\rho)$.
Proof.
The symmetry of the subdivision $\mathcal{R}_{n}^{0}(\rho)$ follows immediately from the symmetry of the metric $\rho$. Moreover as the height of the origin is chosen to be 00, the condition (4.1) is satisfied for $a_{j}=0$ with an equality for any choice $c$. Hence, every facet of $\mathcal{R}_{n}^{0}(\rho)$ contains the origin.
Now fix $I\subseteq[n]^{2}$. The collection $\{p_{ij}\mid(i,j)\in I\}$ is the set of all $p_{ij}$ contained in a facet of $KRW(\rho)$ if and only if
there exists a vector $\textrm{there exists a vector }c\in\mathbb{R}^{n}\textrm{ with }\begin{cases}c\cdot p_{ij}=1&\textrm{ if }(i,j)\in I\\ c\cdot p_{ij}<1&\textrm{ if }(i,j)\not\in I.\end{cases}$
This is equivalent to requiring that
$\begin{cases}c\cdot(e_{i}-e_{j})=\rho_{ij}&\textrm{ if }(i,j)\in I\\ c\cdot(e_{i}-e_{j})<\rho_{ij}&\textrm{ if }(i,j)\not\in I.\end{cases}$
So by definition of $\mathcal{R}_{n}^{0}(\rho)$ (see Remark 3.10) the latter condition is equivalent to the statement that the set $\{e_{i}-e_{j}\mid(i,j)\in I\}\cup\{0\}$ is a maximal cell of the regular subdivision $\mathcal{R}_{n}^{0}(\rho)$. ∎
A fruitful construction of Gel’fand, Kapranov, and Zelevinsky encodes the various regular triangulations of a point configuration $\mathcal{P}$ as points in a polyhedral fan called the secondary fan of $\mathcal{P}$.
Theorem 4.4 ([GKZ, DLRS]).
Let $\mathcal{P}=\{a_{1},\dots,a_{k}\}$ be set of points in $\mathbb{R}^{n}$.
  1. (1)
    The secondary fan $\Sigma(\mathcal{P})$ is a polyhedral fan in $\mathbb{R}^{k}$ such that two points $\omega,\omega^{\prime}\in\mathbb{R}^{k}$ lie in the same open (secondary) cone of $\Sigma(\mathcal{P})$ if and only if the regular subdivisions $\mathcal{P}(\omega)$ and $\mathcal{P}(\omega^{\prime})$ are equal.
  2. (2)
    The fan $\Sigma(\mathcal{P})$ is complete, that is the support of $\Sigma(\mathcal{P})$ is $\mathbb{R}^{k}$, as all points define some regular triangulation.
  3. (3)
    The subdivision $\mathcal{P}(\omega)$ is a refinement of $\mathcal{P}(\omega^{\prime})$ if and only if $\omega^{\prime}$ is in the closure of the open cone that contains $\omega$.
  4. (4)
    The regular triangulations of $\mathcal{P}$ correspond to the top-dimensional open cones of  $\Sigma(\mathcal{P})$.
Remark 4.5.
Proposition 4.3 is the starting point for the study of KRW polytopes through symmetric subdivisions of $\mathcal{R}_{n}^{0}$ where each maximal cell contains the origin. In particular, one can enumerate the combinatorial types of KRW polytopes by enumerating the combinatorial types of such triangulations via the secondary fan of $\mathcal{R}_{n}^{0}$ described above in Theorem 4.4. This approach leads to the enumeration of the combinatorial types of KRW polytopes for the case $n=6$ using the TOPCOM software package [Rambau, Ram23], see Section 8.
Building on Proposition 4.3 we can reconstruct the Wasserstein fan as a ‘symmetrized version’ of the secondary fan of the root polytope:
Proposition 4.6.
Let $\mathcal{R}_{n}^{0}$ be as above. Then, $\Sigma(\mathcal{R}_{n}^{0})$ is a fan in $\mathbb{R}^{n(n-1)+1}$, say with coordinates $x_{ij}$ for the vertices $e_{i}-e_{j}$ for $i\neq j\in\left[n\right]$ and $x_{0}$ for the origin. Then
$\mathscr{F}_{n}=\pi\big{(}\Sigma(\mathcal{R}_{n}^{0})\cap Q\big{)}\cap\mathbf{M}_{n},$
where $\pi:\mathbb{R}^{n(n-1)+1}\to\mathbb{R}^{\binom{n}{2}}$ is the projection sending $x_{ij}$ to $x_{\{i,j\}}$ for $1\leq i<j\leq n$, and $Q:=\{x\in\mathbb{R}^{n(n-1)+1}\mid x_{ij}=x_{ji}\textrm{ for }1\leq i<j\leq n\textrm{ and }x_{0}=0\}$.
Moreover, the above equality is an equality of fans.
Proof.
Let $\rho$ be an $n$-metric in $\mathbf{M}_{n}$ and let $F_{\rho}$ be the unique open face of the Wasserstein fan $\mathscr{F}_{n}$ that contains $\rho$. Furthermore, let $\rho^{\prime}$ be another $n$-metric. By construction of the regular subdivisions (Proposition 4.3.(1)) and by symmetry of $\rho$ and $\rho^{\prime}$, surely $\mathcal{R}_{n}^{0}(\rho)$ and $\mathcal{R}_{n}^{0}(\rho^{\prime})$ lie in $Q$.
Now, by construction of $\mathscr{F}_{n}$ and Theorem 2.5, the polytopes $KRW(\rho)$ and $KRW(\rho^{\prime})$ determine the same collection of admissible graphs and therefore have the same combinatorial structure if and only if $\rho^{\prime}\in F_{\rho}$. By Proposition 4.3, this means that $\mathcal{R}_{n}^{0}(\rho)$ and $\mathcal{R}_{n}^{0}(\rho^{\prime})$ define the same regular subdivision of $\mathcal{R}_{n}^{0}$ if and only if $\rho^{\prime}\in F_{\rho}$. Hence $\rho$ and $\rho^{\prime}$ lie in the same open cone of $\Sigma(\mathcal{R}_{n}^{0})\cap Q$ if and only if $\rho^{\prime}\in F_{\rho}$ which implies the claim. ∎
5. Relationship to tight spans and injective hulls
We now want to relate the stratification of the metric cone by combinatorial type of KRW polytopes with the one in terms of the combinatorial type of a well-known polyhedral complex associated to metric spaces: the injective hull defined by Isbell [Isbell] and rediscovered by Dress [Dress] with the name tight span. The gist of this section is that for $n\geq 5$ the classification of metric spaces by (labeled, resp. unlabeled) combinatorial type of KRW-polytopes is neither finer nor coarser than the classification by (labeled, resp. unlabeled) combinatorial type of tight spans. This incomparability statement remains true when restricting to a ”pruned” version of the tight span, and it implies incomparability of the fan $\mathscr{F}_{n}$ and the secondary fan of the second hypersimplex.
In order to make precise statements, we start by outlining the construction and some key facts, and refer to [Lan13] for a more detailed account and for a thorough analysis of the structure of injective hulls of general metric spaces.
Given an $n$-metric $\rho$ we consider the (unbounded) polyhedron
$P(\rho):=\{x\in\mathbb{R}^{n}\mid x_{i}+x_{j}\geq\rho_{ij}\textrm{ for all }i,j\in[n]\}.$
For $x,y\in\mathbb{R}^{n}$ write $x\leq y$ if and only if $x_{i}\leq y_{i}$ for all $i=1,\ldots,n$ and define
$E(\rho):=\{y\in P(\rho)\mid\textrm{if }x\in P(\rho)\textrm{ and }x\leq y,\textrm{ then }x=y\},$
the set of coordinate-wise minimal elements of $P(\rho)$.
Remark 5.1.
Equivalently, $E(\rho)$ is the set of bounded faces of the polyhedron $P(\rho)$. This fact is stated in [Dress89, Lemma 1]. We refer to [BasicPhyl, Lemma 5.10] for a proof. In particular, the set $E(\rho)$ is naturally a polyhedral complex.
Remark 5.2 (Injective hulls).
A metric space $X^{\prime}$ is called an injective hull of $X$ if $X^{\prime}$ is injective (see, e.g., [Lan13] for a definition) and there is an isometric embedding $h$ of $X$ into $X^{\prime}$ such that every isometric embedding of $X$ into an injective metric space $Y$ factors through $h$. By work of Isbell [Isbell, Theorem 2.1] every metric space has an injective hull that is unique up to isometry.
Isbell’s existence proof in [Isbell, §2] constructs a special realization of the injective hull as a polyhedral complex which, as is detailed in [Lan13], specializes to $E(\rho)$ for instance when $\rho$ is finite. The same construction was independently rediscovered by Dress in [Dress], under the name tight span. Both names have been widely used in the literature and we will use them interchangeably in order to refer to the object we define next.
Definition 5.3 ([Isbell, §2], [Lan13, §3], [Dress, Theorem 3.(v)]).
The tight span of $\rho$ is the realization of the injective hull of $\rho$ given by the set $E(\rho)$ endowed with the metric induced by the supremum-norm of $\mathbb{R}^{n}$, i.e., the distance between $x,y\in E(\rho)$ is $\sup\{|x_{i}-y_{i}|\mid i\in[n]\}$.
Remark 5.4 (Distinguished vertices of $E(\rho)$).
For every $k\in[n]$, the point $h^{(k)}$ defined as $h^{(k)}_{i}=\rho_{ik}$ is a vertex of $E(\rho)$. It is the image of $k$ under the canonical isometric embedding into $E(\rho)$ (see [Dress, Item (1.4)] and Appendix A).
Example 5.5.
Figures 4 and 5 show a schematic view of the complexes $E(\rho)$ associated to the metrics given in Example 6.13. The numbers along the edges give the distance between the edge’s endpoints with respect to the metric of the injective hull. The solid dots are the image of the canonical embedding of the metric space, and are labeled by the corresponding element of $[n]$.
Remark 5.6 (Pruned Tight Span).
We will have use for a subcomplex $E^{\sharp}(\rho)$ of $E(\rho)$ that we suggest to call ”pruned Tight Span” of $\rho$. We give a definition in Definition A.1, and in Theorem A.3 we prove that it can be characterized as being obtained by removing from $E(\rho)$ all vertices of the form $h^{(k)}$ that are contained in a single one-dimensional cell of $E(\rho)$.
The complexes $E^{\sharp}(\rho_{1})$, $E^{\sharp}(\rho_{2})$ for the metrics in Example 6.13 are obtained from the complexes depicted in Figures 4 and 5 by removing all solid-color vertices and the edges they are incident with.
A note of caution: in the computational geometry package polymake [polymake], the tight span $E(\rho)$ is computed by the method metric_extended_tight_span and our “pruned tight span” $E^{\sharp}(\rho)$ by the method metric_tight_span.
Remark 5.7 (Combinatorial types of tight spans).
We will speak of the labeled combinatorial type of tight spans according to the terminology of Example 2.11 with respect to the following labeling of vertices. Some vertices of $E(\rho)$ have a natural labeling by elements of $[n]$ discussed in Remark 5.4. When passing to $E^{\sharp}(\rho)$ via the characterization given in Remark 5.6, for each $k$ such that $h^{(k)}$ is pruned, there is a unique vertex of $E^{\sharp}(\rho)$, called $v^{(k)}$ in Theorem A.3, that is incident to the pruned edge. Note that if $\rho$ is a metric, for different $k$ all $h^{(k)}$ are distinct, but some $v^{(k)}$ may concide. Thus we can label vertices of $E^{\sharp}(\rho)$ either by the label they inherit from $E(\rho)$ or by the set of labels of adjacent pruned vertices in $E(\rho)$.
Now recall the second hypersimplex $\Delta_{2,n}:=\operatorname{conv}\{e_{i}+e_{j}\mid i,j\in[n],\,i\neq j\}\subseteq\mathbb{R}^{n}$. Following the setup of [HJ08], every $\rho\in\mathbb{R}^{{n\choose 2}}$ defines a lift of $\Delta_{2,n}$:
$L(\rho):=\operatorname{conv}\{(e_{i}+e_{j},-\rho_{ij})\mid i,j\in[n],i\neq j\}+\mathbb{R}_{\geq 0}e_{n+1}\subseteq\mathbb{R}^{n+1}.$
The projection of the bounded faces of $L(\rho)$ onto $\mathbb{R}^{n}$ defines the regular subdivision $\Delta_{2,n}(-\rho)$. The stratification of ($\mathbb{R}^{{n\choose 2}}$ according to combinatorial type of this regular subdivision gives rise to the secondary fan $\Sigma(\Delta_{2,n})$, a well-studied object in its own right (see, e.g., [DST]).
Restricting to the metric, resp. pseudometric cone, we obtain the set of strata
$\mathscr{S}_{n}=\{\sigma\cap\mathbf{M}_{n}\mid\sigma\in\Sigma(\Delta_{2,n})\}\quad\quad\overline{\mathscr{S}}_{n}=\{\sigma\cap\overline{\mathbf{M}}_{n}\mid\sigma\in\Sigma(\Delta_{2,n})\}.$
Note that every stratum is a cone, but these are not polyhedral fans.
Remark 5.8.
A well-known fact going back to work of Herrmann and Joswig [HJ08] states that for any two pseudometrics $\rho_{1},\rho_{2}\in\overline{\mathbf{M}}_{n}$ the complexes $E^{\sharp}(\rho_{1})$ and $E^{\sharp}(\rho_{2})$ have the same labeled combinatorial type if and only if they are in the same stratum of $\overline{\mathscr{S}}_{n}$.
Definition 5.9 (Metric fan).
Let ¯$\overline{\mathscr{M}}_{n}$ be the polyhedral fan consisting of all faces of the polyhedral cone $\overline{\mathbf{M}}_{n}$ of all pseudometrics. The metric fan as defined in [CJK24] is the fan
$\operatorname{MF}(n)=\{\mu\cap\sigma\mid\mu\in\overline{\mathscr{M}}_{n},\sigma\in\Sigma(\Delta_{2,n})\}.$
Remark 5.10.
The fan $\operatorname{MF}(n)$ is polyhedral, and its rays are described in detail in [CJK24, Proposition 25]. Note that since all nontrivial faces of ¯$\overline{\mathscr{M}}_{n}$ are contained in the boundary of $\mathbf{M}_{n}$, the stratification given by $\operatorname{MF}(n)$ and $\mathscr{S}_{n}$ coincide in the topological interior $\mathring{\mathbf{M}}_{n}$ of $\mathbf{M}_{n}$. Precisely, we have $\{\sigma\cap\mathring{\mathbf{M}}_{n}\mid\sigma\in\operatorname{MF}(n)\}=\{\tau\cap\mathring{\mathbf{M}}_{n}\mid\tau\in\mathscr{S}_{n}\}.$ Note also that the labeled combinatorial type of $E(\rho)$ is determined by the cone of $\operatorname{MF}(n)$ containing $\rho$ (see Corollary A.4).
The main theorem of this section is the following incomparability result between the Wasserstein fan and the secondary fan of the second hypersimplex.
Theorem 5.11.
The fan $\mathscr{F}_{n}$ is neither a coarsening nor a refinement of the subdivision $\mathscr{S}_{n}$ of the metric cone induced by secondary fan of $\Delta_{2,n}$ for $n\geq 5$. Similarly, the fan ¯$\overline{\mathscr{F}}_{n}$ is neither a coarsening nor a refinement of the metric fan $\operatorname{MF}(n)$ for $n\geq 5$.
Remark 5.12 (The case $n=4$).
The fan $\mathscr{F}_{4}$ is a refinement of the subdivision $\mathscr{S}_{4}$ of the metric cone induced by the secondary fan of $\Delta_{2,4}$. See Figure 8 and Section 8.1.
Theorem 5.11 is a consequence of the following statement.
Proposition 5.13.
For every $n\geq 4$ there are strict metrics for which the injective hulls have the same labeled combinatorial type, but whose KRW-polytopes have different unlabeled combinatorial type. For $n\geq 5$ there are strict metrics for which the KRW-polytopes have the same labeled combinatorial type, but the injective hulls have different unlabeled combinatorial type.
Proof of Theorem 5.11 under assumption of Proposition 5.13.
All metrics cited in Proposition 5.13 are strict, hence they lie in the topological interior of the metric cone and so Remark 5.10 applies. This means that the statement of Proposition 5.13 remains true substituting the injective hulls with the pruned tight span of Remark 5.6. Now the first claim holds by the result of Herrmann and Joswig cited in Remark 5.8, and the second claim follows from Remark 5.10. ∎
We now turn to the proof of Proposition 5.13, for which will need two general facts that we state as lemmas as they may be of independent interest. Recall that if $P$ and $Q$ are two polytopes in $\mathbb{R}^{n}$ both containing the origin and for which the affine hull of $P$ intersects the affine hull of $Q$ only in the origin, then the free sum of $P$ and $Q$ is $P\oplus Q:=\operatorname{conv}(P\cup Q)$, see for instance [Kal88].
Lemma 5.14.
Let $\rho_{1}$, $\rho_{2}$ be metrics on the sets $[n]$ and $\{n,\ldots,n+m\}$, respectively. Consider the function $\rho_{1}\oplus\rho_{2}:[n+m]^{2}\to\mathbb{R}$ defined via
$(\rho_{1}\oplus\rho_{2})(i,j):=\begin{cases}(\rho_{1})_{ij}&\text{if}\quad i,j\leq n,\\ (\rho_{1})_{in}+(\rho_{2})_{nj}&\text{if}\quad i\leq n<j,\\ (\rho_{2})_{ij}&\text{if}\quad n\leq i,j.\end{cases}$
Then $\rho_{1}\oplus\rho_{2}$ is a metric, and we have a direct (free) sum decomposition
$KRW(\rho_{1}\oplus\rho_{2})=KRW(\rho_{1})\oplus KRW(\rho_{2}).$
Proof.
A straightforward check of the definitions shows that $\rho_{1}\oplus\rho_{2}$ is a metric.
The set $X_{1}$ of pairs of distinct elements from $[n]$ is disjoint from the set $X_{2}$ of pairs of distinct elements of $\{n,\ldots,n+m\}$. The KRW polytopes of $\rho_{1}$ and $\rho_{2}$ both contain the origin and lie in the subspaces $\mathbb{R}^{X_{1}}$ and $\mathbb{R}^{X_{2}}$ of $\mathbb{R}^{\binom{n+m}{2}}$, respectively. Since the $X_{i}$ are disjoint, $\mathbb{R}^{X_{1}}\cap\mathbb{R}^{X_{2}}=\{0\}$. Then, by definition $KRW(\rho_{1})\oplus KRW(\rho_{2})=\operatorname{conv}(KRW(\rho_{1})\cup KRW(\rho_{2})).$ For brevity call $K$ the latter polytope, and let $\hat{\rho}:=\rho_{1}\oplus\rho_{2}$. We have to prove that $K=KRW(\hat{\rho})$.
By definition of $\hat{\rho}$, both $KRW(\rho_{1})$ and $KRW(\rho_{2})$ are contained in $KRW(\hat{\rho})$. Therefore, $K$ is contained in $KRW(\hat{\rho})$. The only potential vertices of $KRW(\hat{\rho})$ not contained in $K$ are given by vectors $\hat{p_{ij}}:=\frac{e_{i}-e_{j}}{\hat{\rho}_{ij}}$ with $i\leq n<j$. But by definition of $\hat{\rho}$ we have
$\frac{e_{i}-e_{j}}{\hat{\rho}_{ij}}=\lambda\frac{e_{i}-e_{n}}{{\rho_{1}}_{in}}+\mu\frac{e_{n}-e_{j}}{{\rho_{2}}_{nj}}$
with $\lambda=\frac{({\rho_{1}})_{in}}{(\rho_{1})_{in}+(\rho_{2})_{nj}}$ and $\mu=\frac{({\rho_{2}})_{nj}}{(\rho_{1})_{in}+(\rho_{2})_{nj}}$ being positive numbers with $\lambda+\mu=1$. Therefore $\hat{p_{ij}}$ is in $K$ for all $i\leq n<j$, thus $KRW(\hat{\rho})=K$. ∎
Lemma 5.15.
If $\rho$ is a metric on the set $[n]$ and $t$ a metric on the set $\{n,n+1\}$, then the tight span of $\rho\oplus t$ is obtained from the tight span of $\rho$ by gluing a new edge at the vertex given by $h^{(n)}$ (defined in Remark 5.4).
Proof.
The statement follows from [Miesch, Theorem 1.1] (see also [HJ08, Lemma 2.2]), but we give a self-contained proof here. Write $\lambda:=t_{n\,n+1}=(\rho\oplus t)_{n\,n+1}$, and consider the pseudometrics $\sigma$ and $\delta$ on $[n+1]$ where $\sigma$ is given by $\sigma_{ij}:=\rho_{ij}$ for $j\neq n+1$, $\sigma_{i\,n+1}:=\rho_{in}$, and $\delta:=\lambda\cdot\delta_{[n],\{n+1\}}$ is a (nonzero) multiple of the split pseudometric defined in Example 2.4. As in [BandeltDress], we can apply the construction of the tight span also to pseudometrics, obtaining
$E(\sigma)$ =$=\{x\in\mathbb{R}^{[n+1]}\mid(x_{1},\ldots,x_{n})\in E(\rho),x_{n+1}=x_{n}\},$
$E(\delta)$ =$=\{x\in\mathbb{R}^{[n+1]}\mid x_{1}=x_{2}=\ldots=x_{n},(x_{n},x_{n+1})\in E(t)\}$
=$=\{(\alpha,\ldots,\alpha,\lambda-\alpha)\in\mathbb{R}^{n+1}\mid 0\leq\alpha\leq\lambda\},$
which are polyhedral complexes isomorphic to $E(\rho)$ and $E(t)$, respectively. Let
$\hat{h}^{(n)}:=(\rho_{in},\ldots,\rho_{nn},\rho_{nn})$
denote the image of $h^{(n)}\in E(\rho)$ in $E(\sigma)$, and let [^1] Sum of subsets of Euclidean space is intended element-wise, i.e., as Minkowski sum.
$\hat{E}(\sigma):=E(\sigma)+(0,0,\ldots,0,\lambda),\quad\hat{E}(\delta):=E(\delta)+\hat{h}^{(n)}.$
Since $\hat{E}(\sigma)\cap\hat{E}(\delta)=\{\hat{h}^{(n)}\}$, it will be enough to show that
(5.1) $E(\rho\oplus t)=\hat{E}(\sigma)\cup\hat{E}(\delta),$
since $\hat{E}(\delta)$ (isometric to the unit interval) is joined to $\hat{E}(\sigma)$ (isomorphic to $E(\rho)$) at the unique common point $\hat{h}^{(n)}+(0,\ldots,0,\lambda)$, the image of $h^{(n)}$ in $E(\rho)$.
The inclusion $E(\rho\oplus t)\supseteq\hat{E}(\sigma)\cup\hat{E}(\delta)$ can either be checked directly with Definition 5.3 or deduced from the fact that $\hat{E}(\sigma)$, resp. $\hat{E}(\delta)$ are the images of $E(\rho)$, resp. $E(t)$, under the canonical inclusion into $E(\rho\oplus t)$ described in [Dress, Item (1.11)] [^2] In the language of [Dress, Item (1.11)], we are considering the inclusions of tight spans $\tau_{[n]}$, resp. $\tau_{n+1}$, induced by the extensions of $\rho$ to $\rho\oplus t$, resp. of $t$ to $\rho\oplus t$..
For the other inclusion start by noticing that, obviously, $\rho\oplus t=\sigma+\delta$. The isolation index of $n+1$ with respect to $\rho\oplus t$ (see Definition 6.7 below for the definition) is $\alpha_{[n],\{n+1\}}^{\rho\oplus t}=\lambda$. Therefore [BandeltDress, Theorem 7] applies and shows $P(\rho\oplus t)=P(\sigma)+P(\delta)$. As is explained, e.g., in [BandeltDress, p. 89] [^3] In [BandeltDress] our $E(\rho)$ is denoted by $T(\rho)$., this implies $E(\rho\oplus t)\subseteq E(\sigma)+E(\delta)$. Now for the inclusion $E(\rho\oplus t)\subseteq\hat{E}(\sigma)\cup\hat{E}(\delta)$ it is enough to show that
(5.2) $E(\rho\oplus t)\cap(E(\sigma)+E(\delta))\subseteq\hat{E}(\sigma)\cup\hat{E}(\delta).$
Take $p\in(E(\sigma)+E(\delta))$. In our setup, without loss of generality we can write
[…]
We surely have that $p\in P(\rho\oplus t)$, since:
  • (i)
    for $i,j\in[n]$: $p_{i}+p_{j}=y_{i}+y_{j}+2\alpha\geq\rho_{ij}+2\alpha\geq(\rho\oplus t)_{ij}+2\alpha\geq(\rho\oplus t)_{ij}$;
  • (ii)
    for $i\in[n]$: $p_{i}+p_{n+1}=y_{i}+\alpha+y_{n}+\lambda-\alpha=y_{i}+y_{n}+\lambda\geq\rho_{in}+t_{nn+1}=(\rho\oplus t)_{i,n+1}$.
If $\alpha=0$, then $y\in E(\rho)$ implies that $(p_{1},\ldots,p_{n})$ is already coordinate-wise minimal satisfying (i). In particular, there is $i_{0}\leq n$ with $y_{i_{0}}+y_{n}=\rho_{{i_{0}}n}$. This implies that $p_{n+1}$ cannot be made smaller either, because $p_{{i_{0}}}+p_{n+1}=y_{i_{0}}+y_{n}+\lambda=\rho_{{i_{0}}n}+t_{nn+1}=(\rho+t)_{{i_{0}}n+1}$ is an equality. In this case, then, $p\in E(\rho\oplus t)$ implies $p\in\hat{E}(\sigma)$.
If $\alpha>0$, then $y\in E(\rho)$ implies that none of the inequalities in (i) is an equality, indeed $p_{i}+p_{j}\geq\rho_{ij}+2\alpha$ for all $i,j\in[n]$. Now if $y_{j}+y_{n}>\rho_{in}$ for some $j$, define $\tilde{p}$ by $\tilde{p}_{j}:=p_{j}-{\alpha}$ and $\tilde{p}_{i}:=p_{i}$ for $i\neq j$. Then $\widetilde{p}\in P(\rho\oplus t)$, yet $\tilde{p}$ is smaller than $p$ in the $j$-th coordinate. Thus $p\in E(\rho\oplus t)$ implies that $y_{j}+y_{n}=\rho_{in}$ for all $j\in[n]$, so $p\in\hat{E}(\delta)$. This concludes the proof of Equation 5.2, hence of Equation 5.1 and, with it, of the claim. ∎
Definition 5.16.
For given $k>0$ let $t^{(n,k)}$ denote the $k$-metric defined on the set $\{n,\ldots,n+k\}$ by setting $t^{(k)}_{i,j}:=|j-i|$. This is a graph metric realized on the full set of vertices of a path with $k+1$ vertices.
Remark 5.17.
Let $\rho$ be any $n$-metric. Then
$\rho\oplus t^{(n,k)}=(\cdots(\rho\oplus t^{(n,1)})\oplus t^{(n+1,k)})\cdots)\oplus t^{(n+k-1,1)}.$
By repeated application of Lemma 5.15 we obtain that the tight span of $\rho\oplus t^{(n,k)}$ is obtained from the tight span of $\rho$ by appending a $k+1$-path at the vertex $h_{n}$ of $E(\rho)$.
Proof of Proposition 5.13.
Example 6.13 exhibits two $5$-metrics that lie in the same open cone of  $\mathcal{W}_{5}$ (thus they have combinatorially isomorphic KRW polytopes), but lie in different cones of the secondary fan of $\Delta_{2,5}$ (and thus have different tight spans). Moreover, Example 5.18 shows two metrics on $4$ points that are in the same cone of the secondary fan of $\Delta_{2,4}$ but have non-isomorphic KRW polytopes. Starting from those, we are going to construct examples for any (finite) number of points.
Fix $k>0$ and recall the metrics $t^{(n,k)}$ defined in Definition 5.16. Since these are graph metrics realized on the full set of vertices of a tree with $k+1$ vertices, $KRW(t^{(n,k)})$ is unimodularly equivalent to the cross-polytope of dimension $k$. Moreover, the tight span of $t^{(k)}$ is combinatorially equivalent to the tree realizing it, i.e., a path with $k$ edges, with vertices labeled $n,\ldots,n+k$ (see, e.g., [Dress, Theorem 8]).
Now given an $n$-metric $\rho$ consider the $n+k$-metric $\rho^{(k)}:=\rho\oplus t^{(n,k)}$. By Remark 5.17, the tight span $E(\rho^{(k)})$ is obtained from that of $\rho$ by appending a $k$-path at the vertex labeled $n$. Moreover, by Lemma 5.14 there is a direct (free) sum decomposition $KRW(\rho^{(k)})=KRW(\rho)\oplus KRW(t^{(k)})$, implying that the poset of non-maximal faces of $KRW(\rho^{(k)})$ is the Cartesian product of the posets of non-maximal faces of $KRW(\rho)$ and of $KRW(t^{(k)})$ (see, e.g., [Kal88]).
Let $\rho_{1}$, $\rho_{2}$ be the two $5$-metrics of Example 6.13. The isomorphism between the polytopes $KRW(\rho_{1})$ and $KRW(\rho_{2})$ induces an isomorphism between $KRW(\rho_{1}^{(k)})$ and $KRW(\rho_{2}^{(k)})$. On the other hand, any isomorphism between the tight spans of $\rho_{1}^{(k)}$ and $\rho_{2}^{(k)}$ must map the $t^{(n,k)}$-part to itself (since the tight spans of $\rho_{1}$ and $\rho_{2}$ do not contain any path of length more than $1$ consisting of vertices with degree $\leq 2$, see Example 6.13), and thus would induce an isomorphism between the tight spans of $\rho_{1}$ and $\rho_{2}$. The latter cannot exist, and so we show that $\mathscr{F}_{n}$ is not a refinement of the secondary fan of $\Delta_{2,n}$ for all $n\geq 5$.
Let $\rho_{1}$, $\rho_{2}$ be the two $4$-metrics of Example 5.18. The combinatorial isomorphism between the tight span of $\rho_{1}$ and $\rho_{2}$ extends to an isomorphism of the tight spans of $\rho_{1}^{(k)}$ and $\rho_{2}^{(k)}$. On the other hand the fact that $KRW(\rho_{2})$ is simplicial implies simpliciality of $KRW(\rho_{2}^{(k)})$, and since $KRW(\rho_{1})$ is not simplicial, neither is $KRW(\rho_{1}^{(k)})$. In particular, $KRW(\rho_{1}^{(k)})$ is not isomorphic to $KRW(\rho_{2}^{(k)})$ for any $k>0$. We have proved that for all $n\geq 5$ the fan $\mathscr{F}_{n}$ is not a coarsening of the secondary fan of $\Delta_{2,n}$. ∎
Example 5.18.
Consider the following two $4$-metrics.
$\rho_{1}:=\begin{pmatrix}0&3&3&4\\ 3&0&4&3\\ 3&4&0&3\\ 4&3&3&0\end{pmatrix}\quad\quad\rho_{2}:=\begin{pmatrix}0&4&3&5\\ 4&0&5&3\\ 3&5&0&4\\ 5&3&4&0\end{pmatrix}$
These metrics have combinatorially isomorphic tight spans, but non-isomorphic KRW polytopes, see Figure 2. The polytopes are not combinatorially isomorphic as $KRW(d_{1})$ has two square faces whereas $KRW(d_{2})$ is simplicial.
((a)) $KRW(d_{1})$
((b)) tight span of both metrics
((c)) $KRW(d_{2})$
Figure 2. Illustration for Example 5.18. (a), (c): KRW polytopes of the metrics $\rho_{1}$, $\rho_{2}$. (b): Combinatorial type of the tight span. The horizontal edges are of length $1$ and $2$ for the metrics $\rho_{1}$ and $\rho_{2}$, respectively. All other edges are of length $1$.
6. Noteworthy subfans
6.1. Tree-like metrics
An $n$-metric $\rho$ is called tree-like if it can be represented as the shortest-path metric among a subset of $n$ vertices of a weighted tree $T(\rho)$. It is known since work of Buneman [Buneman] that an $n$-metric $\rho$ is tree-like if and only if it satisfies the four-point condition
(6.1) $\rho_{ij}+\rho_{kl}\leq\max\{\rho_{ik}+\rho_{jl},\rho_{il}+\rho_{jk}\}\textrm{for all }i,j,k,l\in[n].$
These are exactly the metric spaces whose tight span is a tree (i.e., a one-dimensional contractible polyhedral complex), see e.g., [Dress, Theorem 8].
Proposition 6.1.
The set of tree-like metrics is a subfan of $\mathscr{F}_{n}$. In particular, the set of phylogenetic trees on $n$ leaves is a union of (relatively open) cells of $\mathscr{F}_{n}$.
Proof.
A metric $\rho$ satisfies the four-point condition (6.1) for a quadruple $A=\{i,j,k,l\}\in{n\choose 4}$ if and only if, of the three numbers $\rho_{ij}+\rho_{kl},\rho_{ik}+\rho_{jl},\rho_{il}+\rho_{jk}$, two are equal and the third does not exceed the other two. In turn, this is equivalent to saying that there is a permutation $\pi_{A}=(i^{\prime},j^{\prime},k^{\prime},l^{\prime})$ of the elements of $A$ such that
$\rho\in\mathscr{C}_{\pi_{A}}:=H_{(i^{\prime},j^{\prime}),(k^{\prime},l^{\prime})}\cap H_{(i^{\prime},k^{\prime}),(j^{\prime},l^{\prime})}^{+}.$
This means that a metric $\rho$ is tree-like if and only if there is a choice of a permutation $\pi_{A}$ for every $A\in{n\choose 4}$ such that $\rho\in\bigcap_{A\in{n\choose 4}}\mathscr{C}_{\pi_{A}}$. Thus, letting $S_{A}$ denote the group of permutations of the set $A$, the set of tree-like metrics can be expressed as
$\bigcap_{A\in{n\choose 4}}\bigcup_{\pi_{A}\in S_{A}}\mathscr{C}_{\pi_{A}}.$
Since each $\mathscr{C}_{\pi_{A}}$ is a subfan of $\mathscr{F}_{n}$, it follows that the set of all tree-like metrics is a subfan of $\mathscr{F}_{n}$ as well. ∎
6.2. Circular-decomposable metrics
An $n$-metric $\rho$ is circular-decomposable if there is a cyclic ordering $\prec$ of $[n]$ such that
(6.2) $\rho_{ij}+\rho_{kl}\leq\rho_{ik}+\rho_{jl}\textrm{ whenever }i\prec j\prec k\prec l.$
To our knowledge this condition first appeared in work of Kalmanson [kalmanson], and was proved in [CF98] to be equivalent to the ”maximal compatibility” condition studied by Bandelt and Dress in [BandeltDress].
From Equation 6.2 it follows immediately that a point in $\mathbb{R}^{\binom{n}{2}}$ corresponds to a circular-decomposable (pseudo)metric if and only if it is contained in the cone
(6.3) $K_{\prec}:=\bigcap_{i\prec j\prec k\prec l}H^{+}_{(i,l),(j,k)}\textrm{ for some circular ordering }\prec\textrm{ of }[n].$
The following definition is from [Devadoss-petti].
Definition 6.2.
A split $A\uplus B=[n]$ is said to be compatible with a circular ordering $\prec$ if $i\prec j$ for all $i\in A$ and all $j\in B$. Equivalently, we can label the elements of $[n]$ as $x_{1},\ldots,x_{n}$ in such a way that $x_{i}\prec x_{j}$ if $i\in A$ and $j\in B$.
Lemma 6.3.
For every cyclic ordering $\prec$ the set $K_{\prec}$ is a simplicial cone of dimension ${\binom{n}{2}}$. Its rays are generated by the pseudometrics associated to the splits that are compatible with $\prec$. Irrespective of the chosen $\prec$, the cone $K_{\prec}$ contains the cone ¯$\overline{\mathscr{E}}_{n}$ generated by the elementary splits.
Proof.
The last claim follows after noticing that elementary splits are compatible with every circular ordering. The other claims are proved as items (i-iii) in [BandeltDress, Theorem 5, p. 75]. ∎
Theorem 6.4.
The set of all circular decomposable pseudometrics on $n$ points is the full-dimensional simplicial subfan $\overline{\mathscr{K}}_{n}$ of ¯$\overline{\mathscr{F}}_{n}$ given by all points contained in the set
[…]
Accordingly, the circular decomposable metrics form a full-dimensional subfan $\mathscr{K}_{n}$ of $\mathscr{F}_{n}$.
Proof.
Equation 6.3 implies immediately that $\overline{\mathscr{K}}_{n}$ is indeed a subfan of ¯$\overline{\mathscr{F}}_{n}$. The claim about simpliciality and dimension follows from Lemma 6.3. ∎
Remark 6.5 (The space of Circular Split Networks).
A metric is circular decomposable if and only if it can be realized by a circular split network [Devadoss-petti, §2.2]. Devadoss and Petti [Devadoss-petti] define and study a simplicial fan $\operatorname{CSN}_{n}$ that parametrizes all circular split networks on $n$ points associated to circular decomposable metrics without trivial splits.
Corollary 6.6.
The set of circular decomposable metrics without trivial splits defines a subfan of ¯$\overline{\mathscr{F}}_{n}$ consisting of all cones of $\overline{\mathscr{K}}_{n}$ whose closure does not contain any ray of ¯$\overline{\mathscr{E}}_{n}$. This fan structure on the set of circular decomposable metrics without trivial splits is combinatorially isomorphic to the one defined and studied in [Devadoss-petti] and [HKT], see Remark 6.5 above.
6.3. Split decomposition
We now turn to the class of totally split-decomposable spaces, as introduced by Dress and studied by Bandelt and Dress [BandeltDress].
Definition 6.7.
Let $\rho$ be a symmetric function on a set $X$ and let $A,B\subseteq X$. Define the isolation index $\alpha_{A,B}^{\rho}$ as
$\alpha_{A,B}^{\rho}=\frac{1}{2}\cdot\min_{\begin{subarray}{c}a,a^{\prime}\in A\\ b,b^{\prime}\in B\end{subarray}}\left(\max\{a^{\prime}b+b^{\prime}a,a^{\prime}b^{\prime}+ab,aa^{\prime}+bb^{\prime}\}-(aa^{\prime}+bb^{\prime})\right),$
where $uv=\rho(u,v)$.
Remark 6.8.
As noted in [BandeltDress], $\alpha^{\rho}_{A,B}\geq 0$, $\alpha^{\rho}_{A,B}=0$ whenever $A\cap B\not=\emptyset$. Moreover, if $\rho$ is a metric and both $A$ and $B$ have at least 2 elements, then $\alpha^{\rho}_{A,B}=\alpha^{\rho}_{\{a,a^{\prime}\},\{b,b^{\prime}\}}$ for some points $a,a^{\prime}\in A$, $b,b^{\prime}\in B$ with $a\neq a^{\prime}$, $b\neq b^{\prime}$.
In the following, when writing expressions such as $\alpha^{\rho}_{\{a,a^{\prime}\},\{b,b^{\prime}\}}$ we will assume that $a,a^{\prime},b,b^{\prime}$ are pairwise distinct points.
A bipartition $A\uplus B=[n]$ is called a $\rho$-split if $\alpha^{\rho}_{A,B}>0$. A symmetric function $\rho$ which does not admit any $\rho$-splits is called split-prime.
Theorem 6.9 ([BandeltDress], Theorem 2 and Corollary 2).
Every symmetric function $\rho:X\times X\rightarrow\mathbb{R}$ on a finite set $X$ can be expressed as
$\rho=\rho_{0}+\sum\alpha^{\rho}_{A,B}\cdot\delta_{A,B}$
with $\rho_{0}$ split-prime. The metric $\rho_{0}$ is uniquely determined.
Definition 6.10.
A metric $\rho$ is called totally split-decomposable if it can be written as a positively weighted sum of split metrics. Equivalently, $\rho_{0}=0$ in the above expression.
The following is a criterion for checking whether a metric is totally split-decomposable:
Theorem 6.11 ([BandeltDress], Theorem 6).
Let $\rho:X\times X\rightarrow\mathbb{R}$ be a symmetric function with zero diagonal.
The following conditions are equivalent:
  1. (1)
    $\rho$ is a totally split-decomposable metric.
  2. (2)
    For all $t,u,v,w,x\in X$
    $\alpha_{\{t,u\},\{v,w\}}\leq\alpha_{\{t,x\},\{v,w\}}+\alpha_{\{t,u\},\{v,x\}}.$
Remark 6.12.
From Theorem 6.11 it is clear that all metrics on $4$ points are split-decomposable. For $|X|\geq 5$ we can ask the question whether two metrics in the same cone of the Wasserstein arrangement are either both split-decomposable or both not split-decomposable. The next example answers this question in the negative.
Example 6.13.
Consider the two metrics $(\rho_{1},\rho_{2})$ on five points given by the following matrices:
$\rho_{1}=\begin{bmatrix}0&125&48&149&84\\ 125&0&149&48&99\\ 48&149&0&125&77\\ 149&48&125&0&92\\ 84&99&77&92&0\\ \end{bmatrix}~{}~{}~{}\rho_{2}=\begin{bmatrix}0&7447&4316&10083&5584\\ 7447&0&10083&4316&5199\\ 4316&10083&0&7447&7560\\ 10083&4316&7447&0&6179\\ 5584&5199&7560&6179&0\end{bmatrix}$
These metrics lie in the same cone of the Wasserstein arrangement $\mathcal{W}_{5}$, and therefore the associated KRW polytopes have the same combinatorial structure. The KRW polytope of $\rho_{1}$ is depicted in Figure 3. Sketches of the tight spans are depicted in Figures 4 and 5.
The metric $\rho_{1}$ is split-decomposable but for the metric $\rho_{2}$ we find
$\alpha^{\rho_{2}}_{\{2,3\},\{1,5\}}>\alpha^{\rho_{2}}_{\{3,4\},\{1,5\}}+\alpha^{\rho_{2}}_{\{2,3\},\{4,5\}}.$
By Theorem 6.11, this certifies that $\rho_{2}$ is not split-decomposable.
Figure 3. Schlegel diagram of the polar dual of the KRW polytope (i.e., the Lipschitz polytope) of the metric $\rho_{1}$ from Example 6.13.
Figure 4. Sketch of the tight span of the metric $\rho_{1}$ in Example 6.13. The complex $E(\rho_{1})$ has dimension $2$ and consists of four rectangular faces and five dangling edges. It has $14$ total vertices, five of which (the solid dots in the picture) are the image of the canonical embedding of the metric space into its injective hull. We have labeled some edges with a number corresponding to the distance of the edge’s endpoints with respect to the distance in the injective hull, which is induced on $E(\rho_{1})\subseteq\mathbb{R}^{5}$ by the uniform norm (”supremum-norm”). Edges that appear parallel in the picture have the same ”length”. See also Remark 6.14.
Figure 5. Sketch of the tight span of the metric $\rho_{2}$ in Example 6.13. The complex $E(\rho_{2})$ has $5$ two-dimensional cells (two pentagons, two quadrilaterals, one triangle) and five dangling edges. This metric is ”generic of type II” in the sense of Bandelt and Dress [BandeltDress].
Remark 6.14.
Note that the notion of ”generic” metric discussed in [BandeltDress] is not the same as the one of Definition 1.1. Both metrics in Example 6.13 are generic in the sense of our Definition 1.1, but only $\rho_{2}$ is generic in the sense of [BandeltDress], while $\rho_{1}$ is ”degenerate” in the sense of Bandelt and Dress because the metrics used in the canonical decomposition of $\rho_{1}$ are a strict subset of those for $\rho_{2}$ (indeed $\rho_{2}$ has an additional nonzero contribution from a split-prime metric).
Remark 6.15.
Restricting to the metrics of a specific cone of the Wasserstein Arrangement $\mathcal{W}_{5}$, the conditions of Theorem 6.11 form a local hyperplane arrangement within this cone where the split-decomposable metrics lie in exactly one chamber.
Indeed, taking a look at the condition for split-decomposability for $n=5$, since the isolation indices are defined on sets of two elements, the minimum runs just over one possibility. For the maximum, this is determined by the chamber of $\mathcal{W}_{5}$ the metric is in: the maximum over the three values can be computed by pairwise comparison, but comparing the values $ab+a^{\prime}b^{\prime}$ and $a^{\prime}b+ab^{\prime}$ is nothing other than asking which of the two edge pairs $\{(a,b),(a^{\prime},b^{\prime})\},\{(a^{\prime},b),(a,b^{\prime})\}$ is admissible.
6.4. Consistent totally split-decomposable metrics
Let $\rho$ be a totally split-decomposable metric. In [SixPoints, Theorem 1] (see also [AnExplicit]) it is proved that the tight span of $\rho$ is combinatorially isomorphic to the so-called Buneman complex if and only if $\rho$ satisfies the following six-point condition:
[…]
6.4 For all $I\subseteq[n]$ with $|I|=6$ there are distinct elements $i,j\in I$ with $\rho_{ij}+\rho_{kl}\leq\max\{\rho_{ik}+\rho_{jl},\rho_{jk}+\rho_{il}\}$ for all pairs of distinct $k,l\in I\setminus\{i,j\}$.
Proposition 6.16.
The set of metrics that satisfy the six-point condition Section 6.4 is a subfan of ¯$\overline{\mathscr{F}}_{n}$.
Proof.
It is enough to prove that the set of all $\rho\in\mathbb{R}^{{\binom{n}{2}}}$ satisfying the six-point condition is a union of faces of the Wasserstein arrangement. For this, notice that the inequality in Section 6.4 is satisfied whenever $\rho\in T_{ijkl}$, where
$T_{ijkl}=H^{+}_{(i,k),(j,l)}\cup H^{+}_{(i,l),(j,k)}$
is a union of (closed) chambers of $\mathcal{W}_{n}$. Then, $\rho$ satisfies the six-point condition in full if and only if
$\rho\in\bigcap_{\begin{subarray}{c}I\subseteq[n]\\ |I|=6\end{subarray}}\;\;\bigcup_{\begin{subarray}{c}i,j\in I\\ i\neq j\end{subarray}}\;\;\bigcap_{\begin{subarray}{c}k,l\in I\setminus\{i,j\}\\ k\neq l\end{subarray}}T_{ijkl}.$
This proves that the set of $\rho$ that satisfy the six-point condition is obtained uniting and intersecting sets of (closed) chambers, hence it is a union of faces of the arrangement ¯$\overline{\mathscr{F}}_{n}$. ∎
Remark 6.17.
A totally split-decomposable metric that satisfies (6.4) is called consistent in [SixPoints]. We note that the set of consistent totally split-decomposable metrics is not a subfan of the Wasserstein fan.
To prove this, let $\rho_{1}$, $\rho_{2}$ be the $5$-metrics defined in Example 6.13, let $t^{(5,2)}$ the ”path metric on two edges” defined in Definition 5.16, and recall the ”direct sum” construction of Lemma 5.14. The proof of Theorem 5.11 shows that the two $7$-metrics
$\rho_{1}\oplus(12\cdot t^{(5,2)}),\quad\quad\rho_{2}\oplus(840\cdot t^{(5,2)}),$
are in the same cone of the Wasserstein fan $\mathcal{W}_{7}$, and a computer check reveals that both satisfy the six-point condition (6.4). However, the first metric is totally split-decomposable, while the second is not. (This can be either checked by computer or proved noticing that for every $n$-metric $\rho$ the metric $\rho\oplus(\lambda\cdot t^{(n,k)})$ decomposes as the sum of $\rho$ and $k$ elementary splits, and using the fact that $\rho_{1}$ is split-decomposable but $\rho_{2}$ is not, see Example 6.13.) This shows that consistent totally split-decomposable metrics do not form a subfan of ¯$\overline{\mathscr{F}}_{n}$.
7. The $f$-vector in the strict case
The starting point of this section is the following result by Gordon and Petrov. Recall from Definition 1.1 the definition of a generic metric.
Theorem 7.1 ([GordonPetrov], Theorem 1).
Let $X,|X|=n+1$ be a metric space with generic metric $\rho$. Then for $0\leq m\leq n$, the number of $(n-m)$-dimensional faces of the Lipschitz polytope of $\rho$ is
$\binom{n+m}{m,m,n-m}.$
The proof of this theorem uses the admissible graphs discussed earlier in Section 2.2. In fact, using [GordonPetrov, Corollary 1.(3)] the number of faces of dimension $(n-m)$ in the generic case agrees with the number of different admissible (directed) forests with $m$ edges on $n+1$ vertices.
The strict, non-generic case is somewhat more complicated, but general formulae for faces of dimension 00 and $1$ can be stated by using the graphic characterization of the faces: A set of vertices $F$ of the polytope defines a face if the intersection of all facets containing $F$ is identical to $F$. In order to check whether a given (admissible) graph $G$ is the graph $G(F)$ for some a face $F$ of the polytope (see Section 2.2), we need to check that for every edge $e$ not in $G$, there is a maximal admissible graph (i.e., a facet) containing all edges of $G$, but not $e$.
Corollary 7.2 ([GordonPetrov] Corollary 1.(2)).
Let $\rho$ be a strict metric, then the number of vertices satisfy $f_{0}(KRW(\rho))=n\cdot(n+1)$.
For $k\in\mathbb{N}$, denote by
$\mathcal{W}_{n}^{k}:=\left\{H_{a,b}\left|\begin{array}[]{l}a,b\in[n]^{k}\\ a_{1},\ldots,a_{k},b_{1},\ldots,b_{k}\textrm{ mutually distinct}\end{array}\right.\right\}$
the set of hyperplanes in $\mathcal{W}_{n}$ corresponding to cycles of length $2k$ in the complete graph $K_{n}$.
Definition 7.3.
Let $\rho$ be a strict metric on $l$ points and define
$X_{\rho}:=\bigcap_{\begin{subarray}{c}H\in\mathcal{W}_{l}\\ \rho\in H\end{subarray}}H\quad\textrm{ and }\quad r_{k}(\rho):=|\{H\in\mathcal{W}_{l}^{k}\mid X_{\rho}\subseteq H\}|.$
Lemma 7.4.
Let $\rho$ be a strict metric which is not contained in any hyperplane of $\mathcal{W}_{l}^{2}$. Then for each quadruple $x_{1},y_{1},x_{2},y_{2}$ of vertices, if the edge set $\{(x_{1},y_{1}),(x_{2},y_{2})\}$ is admissible, it is face-defining.
Proof.
Let $\{(x_{1},y_{1}),(x_{2},y_{2})\}$ be a pair of admissible edges and let $e=(i,j)$ be an edge which is admissible together with the pair. We need to show that there is an admissible graph containing the pair, but not $e$, such that the addition of $e$ would turn it inadmissible. We use the strictness of the metric and the strict inequality:
(7.1) $\rho(x_{1},y_{1})+\rho(x_{2},y_{2})<\rho(x_{1},y_{2})+\rho(x_{2},y_{1}).$
The trivial case is $|e\cap\{x_{1},y_{1},x_{2},y_{2}\}|=2$, then $e$ is either of $(x_{1},y_{2}),(x_{2},y_{1})$. In this case, the graph we are looking for is the pair of edges $\{(x_{1},y_{1}),(x_{2},y_{2})\}$ and the other edge, i.e., $(x_{2},y_{1})$ or $(x_{1},y_{2})$ since the set of edges $\{(x_{1},y_{1}),(x_{2},y_{2}),(x_{1},y_{2}),(x_{2},y_{1})\}$ is inadmissible by assumption that $\rho$ is contained in no hyperplane in $\mathcal{W}_{l}^{2}$.
For $|e\cap\{x_{1},y_{1},x_{2},y_{2}\}|=1$ or 00: If $i$ is not contained in the quadruple, consider the edges $(x_{1},i),(x_{2},i)$, if $j$ is not contained in the quadruple, consider the edges $(j,y_{1}),(j,y_{2})$. If we prove that one of these edges is admissible, we are finished: due to the strictness of $\rho$, the addition of $e$ would be inadmissible in either case, since $(i,j)$ together with any of these 4 edges would create a directed path.
We outline the argument for the edges $(j,y_{1}),(j,y_{2})$. Assume to the contrary that both are inadmissible, meaning we have strict equations
$\rho(j,y_{2})+\rho(x_{2},y_{1})<\rho(j,y_{1})+\rho(x_{2},y_{2})$
$\rho(j,y_{1})+\rho(x_{1},y_{2})<\rho(j,y_{2})+\rho(x_{1},y_{1}).$
Adding the two equations, we get a contradiction to the inequality (7.1). ∎
With this, we can obtain the following general statement.
Proposition 7.5.
Let $\rho$ be a strict metric on $l=n+1$ points. Then the number of edges satisfies
$f_{1}(KRW(\rho))=\binom{n+2}{2,2,n-2}-2\cdot r_{2}(\rho).$
Note that the multinomial coefficient is the number of faces in the generic case, where $X_{\rho}$ is the empty intersection.
Proof.
Let $\rho$ be a strict metric. By Corollary 2.8, admissible graphs on $2$ edges are of one of the two types depicted in Figure 6.
$x_{1}$$x_{2}$$y_{1}$$y_{2}$
((a)) Type (a)
$x_{1}=x_{2}$$y_{1}$$y_{2}$
$x_{1}=x_{2}$$y_{1}$$y_{2}$
((b)) Type (b)
Figure 6. The two types of admissible graphs on two edges.
Each of the $\frac{(n+1)!}{(n-1)!}$ graphs of type (b) is always trivially admissible. Let us count the admissible graphs of type (a): for every choice of an ordered $4$-tuple $x_{1},y_{1},x_{2},y_{2}$ consider the quantities
$s^{+}:=\rho(x_{1},y_{1})+\rho(x_{2},y_{2}),\quad\quad s^{-}:=\rho(x_{1},y_{2})+\rho(x_{2},y_{1}).$
There are three cases: either $s^{+}<s^{-}$, in which case the graph with the edge set $\{(x_{1},y_{1}),(x_{2},y_{2})\}$ is admissible and the one with the edge set $\{(x_{1},y_{2}),(x_{2},y_{1})\}$ is not, or $s^{+}>s^{-}$, in which case the graph with the edge set $\{(x_{1},y_{2}),(x_{2},y_{1})\}$ is admissible and the one with the edge set $\{(x_{1},y_{1}),(x_{2},y_{2})\}$ is not, or $s^{+}=s^{-}$, in which case $\rho$ is contained in $H_{(x_{1},x_{2}),(y_{1},y_{2})}\in\mathcal{W}_{l}^{2}$.
If $r_{2}(\rho)=0$ then the third case never enters and so exactly one of the other alternatives is realized. Moreover, in this setting the admissible graphs obtained in the first two cases are face-defining according to Lemma 7.4, and we obtain a total of $\frac{(n+1)!}{4(n-3)!}$ admissible graphs of type (a). Thus, when $\rho$ is not contained in any member of $\mathcal{W}_{l}^{2}$ the proposition holds.
If $r_{2}(\rho)>0$, let $H_{1},\ldots,H_{r_{2}(\rho)}$ be the hyperplanes of $\mathcal{W}_{l}^{2}$ containing $\rho$. The count of subgraphs of type (b) does not change. We now consider the graphs of type (a). Let $C_{H}$ be a cycle corresponding to $H\in\mathcal{W}_{l}^{2}$ (see Remark 3.6). What happens if we move from either halfspace onto the hyperplane $H$ is that all graphs remain admissible. Moreover, if an admissible graph $G$ contains the set $C_{H}^{+}$ or $C_{H}^{-}$, respectively, the union of $G$ with $C_{H}^{-}$ or $C_{H}^{+}$, respectively, becomes admissible, too. Thus, $C_{H}^{+},C_{H}^{-}$ are no longer faces of the polytope and we therefore subtract them from the count. Faces of dimension $1$ which correspond to paths on three vertices are not affected.
Without loss of generality, we can regroup the hyperplanes $H_{1},\ldots,H_{r_{2}(\rho)}$ by the ground set of their underlying cycles. Each quadruple $\{u,v,w,t\}$ either appears zero, once or three times, since the intersection
$H_{(v,w),(u,t)}\cap H_{(v,u),(w,t)}\cap H_{(v,t),(w,u)}$
has codimension $2$ for any quadruple $\{u,v,w,t\}$. Thus if $\rho$ is in the intersection of two hyperplanes of cycles of one quadruple, it automatically is in the intersection of all three of them.
The first case is trivial, the case of the ground set appearing once is described above and in the last case, we subtract all $6$ faces ($2$ per cycle) from the generic count. Doing this iteratively for each quadruple gives the result. e therefore subtract them from the count. Faces of dimension $1$ which correspond to paths on three vertices are not affected. ∎
For a low-dimensional example, see Table 2, where the $f$-vectors and corresponding inequalities are listed for $l=4$.
8. Explicit computations
In this section, we explicitly study the KRW polytopes of metrics on $4$, $5$, and $6$ points. We published an accompanying data set of one generic metric of each combinatorial type of these sizes on Zenodo [zenodo].
8.1. The case $n=4$
There is only one generic KRW polytope up to symmetry. The Wasserstein arrangement is depicted in Figure 7 where the cycles corresponding to the hyperplanes are shown next to them.
In this case, the arrangement agrees with the braid arrangement $\mathcal{A}_{3}$ up to the lineality space and thus has 6 chambers.
The strict but not generic cases are illustrated in items (2-4) in Table 2. Items 2 and 3 are metrics on the rank one intersections of the hyperplane arrangement, and item 4 is on a rank 2 intersection. Observe that while the $f$-vector only depends on the rank of the intersection while their tight spans differ.
Figure 7. Wasserstein arrangement for $n=4$. Each hyperplane corresponds to a cycle as described in Remark 3.6 and each chamber has one set of oriented edges of each cycle that is admissible. We obtain the admissible graphs of a neighboring chamber by applying the permutation specified on the hyperplane separating the chambers. The edge pairs in the gray area are the ones corresponding to the horizontal hyperplane.
Nr. case sample metric $f$-vector polytope tight span
1 $d_{12|34}>d_{13|24}>d_{14|23}$ $\begin{pmatrix}0&8&7&5\\ 8&0&5&7\\ 7&5&0&8\\ 5&7&8&0\end{pmatrix}$ (12 30 20) [Uncaptioned image] [Uncaptioned image]
2 $d_{12|34}>d_{13|24}=d_{14|23}$ $\begin{pmatrix}0&6&5&5\\ 6&0&5&5\\ 5&5&0&6\\ 5&5&6&0\end{pmatrix}$ (12 28 18) [Uncaptioned image] [Uncaptioned image]
3 $d_{12|34}=d_{13|24}>d_{14|23}$ $\begin{pmatrix}0&3&3&2\\ 3&0&2&3\\ 3&2&0&3\\ 2&3&3&0\end{pmatrix}$ (12 28 18) [Uncaptioned image] [Uncaptioned image]
4 $d_{12|34}=d_{13|24}=d_{14|23}$ $\begin{pmatrix}0&1&1&1\\ 1&0&1&1\\ 1&1&0&1\\ 1&1&1&0\end{pmatrix}$ (12 24 14) [Uncaptioned image] [Uncaptioned image]
Table 2. Types of strict metrics for $n=4$, their KRW polytopes and the corresponding tight spans.
Figure 8. Wasserstein arrangement for $n=4$, coded by the tight spans. For detailed encoding of the arrangement see Figure 7 and the numbers 1-4 correspond to the tight spans in Table 2. Number 4 is the cone of convex combinations of elementary splits.
8.2. The case $n=5$
The characteristic polynomial of the Wasserstein arrangement $\mathcal{W}_{5}$ is $\chi_{\mathcal{W}_{5}}(t)=t^{10}-15t^{9}+90t^{8}-260t^{7}+350t^{6}-166t^{5}$ which does not factor over the integers. Thus by Zaslavsky’s theorem, the arrangement $\mathcal{W}_{5}$ has $882$ chambers which agrees with the number of maximal cones of the fan $\mathscr{F}_{5}$ [Zas75]. There are 12 generic KRW polytopes up to symmetry that are pairwise not isomorphic. The $f$-vector of these polytopes is the vector (20  90  140  70)209014070(20\,\,90\,\,140\,\,70)( 20 90 140 70 ), as was proved in [GordonPetrov].
Moreover, there are 65 combinatorial types of strict but not generic KRW polytopes. We computed these using the polymake [polymake] hyperplane arrangement package [polymake_arrangements] by enumerating all cells of the Wasserstein arrangement. We record descriptions of these 65 combinatorial types and their metrics in Appendix B.
Now, we proceed with a description of the $12$ combinatorial types of generic KRW polytopes. The encoding of the chambers is to be understood as follows: The even cycles on 5 points are just the 4-cycles. Each 4-cycle can be encoded with the notation as seen in Figure 7 by the cyclic permutations of length two. Consider the following matrix. The $i$-row of this matrix contains the cycles on the ground set $[5]\backslash\{i\}$. For example $M_{12}$ encodes the cycle $[2,3,5,4]$, whereas $M_{22}$ encodes the cycle $[1,3,5,4]$.
$M=\begin{bmatrix}(35)&(45)&(34)\\ (35)&(45)&(34)\\ (25)&(45)&(24)\\ (25)&(35)&(23)\\ (24)&(34)&(23)\end{bmatrix}$
Again using Figure 7, we define the side of the horizontal pair of edges to be positive and can thus encode the structure of the chamber in a $(5\times 3)$-matrix with 1, -1 entries according to the sign of the chamber regarding the hyperplane specified by the above encoding.
The symmetric group $S_{5}$ acts on the chambers by coordinate permutation. The number of chambers yielding the same metric as a given chamber is 120 divided by the order of the stabilizer of the chamber under this action.
In particular, two chambers yield combinatorially equivalent polytopes if and only if there is $\sigma\in S_{5}$ mapping one chamber to the other.
Table 3. Types of generic metrics for $n=5$.
Nr. Chambers sample metric Schlegel diagram stabilizer
1 $\begin{bmatrix}1&1&1\\ 1&1&1\\ 1&1&1\\ 1&1&1\\ 1&1&1\end{bmatrix}$ $\begin{pmatrix}0&16&13&11&9\\ 16&0&9&9&11\\ 13&9&0&9&13\\ 11&9&9&0&16\\ 9&11&13&16&0\end{pmatrix}$ [Uncaptioned image] $\langle(15)(24)\rangle$
2 $\begin{bmatrix}1&1&1\\ 1&1&1\\ 1&1&1\\ 1&1&1\\ 1&-1&1\end{bmatrix}$ $\begin{pmatrix}0&14&14&11&8\\ 14&0&8&9&10\\ 14&8&0&7&12\\ 11&9&7&0&15\\ 8&10&12&15&0\end{pmatrix}$ [Uncaptioned image] id
3 $\begin{bmatrix}1&1&1\\ 1&1&1\\ 1&1&1\\ 1&1&1\\ 1&1&-1\end{bmatrix}$ $\begin{pmatrix}0&48&35&34&26\\ 48&0&31&26&34\\ 35&31&0&26&40\\ 34&26&26&0&48\\ 26&34&40&48&0\end{pmatrix}$ [Uncaptioned image] id
4 $\begin{bmatrix}1&1&1\\ 1&1&1\\ 1&1&1\\ 1&1&1\\ 1&-1&-1\end{bmatrix}$ $\begin{pmatrix}0&12&14&12&7\\ 12&0&10&10&9\\ 14&10&0&6&13\\ 12&10&6&0&15\\ 7&9&13&15&0\end{pmatrix}$ [Uncaptioned image] id
5 $\begin{bmatrix}1&1&1\\ 1&1&1\\ 1&1&1\\ 1&1&1\\ -1&-1&-1\end{bmatrix}$ $\begin{pmatrix}0&12&13&14&6\\ 12&0&13&10&10\\ 13&13&0&7&13\\ 14&10&17&0&16\\ 6&10&13&16&0\end{pmatrix}$ [Uncaptioned image] $\langle(13)(45)\rangle$
6 $\begin{bmatrix}1&1&1\\ 1&1&1\\ 1&1&1\\ 1&1&1\\ -1&1&-1\end{bmatrix}$ $\begin{pmatrix}0&42&35&40&20\\ 42&0&37&26&34\\ 35&37&0&23&37\\ 40&26&23&0&48\\ 20&34&37&48&0\end{pmatrix}$ [Uncaptioned image] id
7 $\begin{bmatrix}1&1&-1\\ 1&1&1\\ 1&1&1\\ 1&1&1\\ 1&1&-1\end{bmatrix}$ $\begin{pmatrix}0&54&36&41&25\\ 54&0&36&25&41\\ 36&36&0&36&36\\ 41&25&36&0&54\\ 25&41&36&54&0\end{pmatrix}$ [Uncaptioned image] $\langle(12)(45),(14)(25)\rangle$
8 $\begin{bmatrix}1&-1&1\\ 1&1&1\\ 1&1&1\\ 1&1&1\\ 1&-1&1\end{bmatrix}$ $\begin{pmatrix}0&48&52&38&23\\ 48&0&26&40&37\\ 52&26&0&26&51\\ 38&40&26&0&49\\ 23&37&51&49&0\end{pmatrix}$ [Uncaptioned image] $\langle(15)(24)\rangle$
9 $\begin{bmatrix}1&-1&1\\ 1&1&1\\ 1&1&1\\ 1&1&1\\ -1&-1&1\end{bmatrix}$ $\begin{pmatrix}0&12&16&13&6\\ 12&0&10&13&10\\ 16&10&0&7&16\\ 13&13&7&0&15\\ 6&10&16&15&0\end{pmatrix}$ [Uncaptioned image] $\langle(14)(35)\rangle$
10 $\begin{bmatrix}1&1&1\\ 1&1&1\\ 1&1&-1\\ 1&1&1\\ 1&1&1\end{bmatrix}$ $\begin{pmatrix}0&48&37&29&31\\ 48&0&29&31&29\\ 37&29&0&29&37\\ 29&31&29&0&48\\ 31&29&37&48&0\end{pmatrix}$ [Uncaptioned image] $\langle(15)(24)\rangle$
11 $\begin{bmatrix}1&1&1\\ 1&1&1\\ 1&1&-1\\ 1&1&1\\ 1&-1&1\end{bmatrix}$ $\begin{pmatrix}0&48&52&35&34\\ 48&0&32&43&26\\ 52&32&0&29&40\\ 35&43&29&0&57\\ 34&26&40&57&0\end{pmatrix}$ [Uncaptioned image] $\langle(23)(45)\rangle$
12 $\begin{bmatrix}1&-1&1\\ 1&1&1\\ 1&1&-1\\ 1&1&1\\ 1&-1&1\end{bmatrix}$ $\begin{pmatrix}0&36&40&23&21\\ 36&0&23&40&21\\ 40&23&0&23&38\\ 23&40&23&0&38\\ 21&21&38&38&0\end{pmatrix}$ [Uncaptioned image] $D_{10}$ in $S_{5}$ ( [^4] We write $D_{10}$ for the dihedral group with $10$ elements, see Table 5.)
8.3. The case $n=6$.
For $6$-point metrics we employed two approaches. First, we computed the number of chambers of the arrangement $\mathcal{W}_{6}$ using the julia module CountingChambers.jl [BEK23]. This is an arrangement in $\mathbb{R}^{15}$ with a $6$-dimensional lineality space consisting of $105$ hyperplanes, see Proposition 3.7. We found that this arrangement has $6,677,863,200$ chambers (this agrees by the above discussion with the number of combinatorial types of generic KRW polytopes).
As explained in Section 4, the generic KRW polytopes correspond exactly to certain regular triangulations of $RP^{0}_{n}$, the (full) root polytope together with the origin, see Proposition 4.3. Jörg Rambau computed all symmetric regular triangulations of $RP^{0}_{6}$ up to symmetry using the latest version of his software TOPCOM [Ram23]. There are exactly $25,224$ such triangulations. This number therefore agrees with the number of combinatorial types of KRW polytopes of $6$-point metrics up to the action of the symmetry group $S_{6}$. We record both statements in the following theorem.
Theorem 8.1.
There are exactly $6,677,863,200$ combinatorial types of generic KRW polytopes corresponding to $6$-point metrics. Under the action of the symmetric group $S_{6}$ these polytopes come in $25,224$ orbits.
[…]
Appendix A Polyhedral complexes associated to pseudometrics
Our goal in this appendix is to describe in an explicit and self-contained way the difference between the tight span $E(\rho)$ and the polyhedral complex $E^{\sharp}(\rho)$ that we call ”pruned tight span” in Section 5 (see Remarks 5.6 and 5.7). The distinction is in order because in general the complexes $E^{\sharp}(\rho)$ and $E(\rho)$ do not coincide, although they are sometimes called by the same name in the literature.
Recall the polyhedron
$P(\rho):=\{x\in\mathbb{R}^{n}\mid x_{i}+x_{j}\geq\rho_{ij}\textrm{ for }i,j\in[n]\}$
and let $E(\rho)$ be the complex of bounded faces of $P(\rho)$. As we have seen in Section 5, $E(\rho)$ is the set underlying a realization of the injective hull of $\rho$ (also named ”tight span” in [Dress]).
Definition A.1.
Given $\rho\in\mathbb{R}^{{n\choose 2}}$, define
$P^{\sharp}(\rho):=\{x\in\mathbb{R}^{n}\mid x_{i}+x_{j}\geq\rho_{ij}\textrm{ for }i,j\in[n],\,i\neq j\}$
and let $E^{\sharp}(\rho)$ be the complex of bounded faces of $P^{\sharp}(\rho)$.
In order to study the relationship between $E^{\sharp}(\rho)$ and $E(\rho)$ we will start by defining, for every $k\in[n]$, two points $h^{(k)},v^{(k)}\in\mathbb{R}^{n}$ as follows:
(A.1) $h^{(k)}_{i}$ $:=d_{ki}\quad\textrm{ for all }i;$
(A.2) $v^{(k)}_{i}$ :=$:=\begin{cases}\xi_{k}&i=k\\ d_{ik}-\xi_{k}&\textrm{ otherwise}\end{cases},\quad\textrm{ where }$
(A.3) $\xi_{k}$ $:=\frac{1}{2}\min\{d_{ik}+d_{kj}-d_{ij}\mid k\not\in\{i,j\}\}.$
Remark A.2.
Note that in the terminology of [BandeltDress] the number $\xi_{k}$ is the isolation index of $k$. For every pseudometric $\rho$ and every $k\in[n]$ we have $\xi_{k}\geq 0$; moreover, $\xi_{k}=0$ if and only if $v^{(k)}=h^{(k)}$.
We will prove the following statement.
Theorem A.3.
The points $v^{(k)}$ for $k\in[n]$ are vertices of the complex $E^{\sharp}(\rho)$, and the latter is a polyhedral subcomplex of $E(\rho)$. More precisely, $E(\rho)$ is obtained by adding to $E^{\sharp}(\rho)$ the vertex $h^{(k)}$ and the edge $e^{(k)}:=\operatorname{conv}\{h^{(k)},v^{(k)}\}$ for every $k$ with $\xi_{k}>0$.
Before proving the theorem we state a corollary about the metric fan $\operatorname{MF}(n)$. Recall the natural labelings of tight spans in Remark 5.7.
Corollary A.4.
The labeled combinatorial type of $E(\rho)$ is determined by the face of $\operatorname{MF}(n)$ containing $\rho$.
Proof.
We know by Remark 5.8 that the labeled combinatorial type of $E^{\sharp}(\rho)$ is determined by the stratum of $\overline{\mathscr{S}}_{n}$ containing $\rho$, and the latter is a union of cones of $\operatorname{MF}(n)$. Since by Theorem A.3 the complex $E(\rho)$ differs from $E^{\sharp}(\rho)$ by attaching one dangling edge at $v^{(k)}$ for every $k$ with $\xi_{k}>0$, it is enough to show that, for every $k\in[n]$, the set of all pseudometrics with $\xi_{k}=0$ is a union of faces of the pseudometric cone $\overline{\mathbf{M}}_{n}$. Indeed, let $K_{ij}^{k}$ be the hyperplane in ($\mathbb{R}^{{n\choose 2}}$ with equation $x_{ik}+x_{kj}-x_{ij}=0$. This hyperplane clearly supports $\overline{\mathbf{M}}_{n}$, hence $F_{ij}^{k}:=\overline{\mathbf{M}}_{n}\cap K_{ij}^{k}$ is a face of $\overline{\mathbf{M}}_{n}$. The set of all pseudometrics with $\xi_{k}=0$ is the union $\bigcup_{i,j\neq k}F_{ij}^{k}$. ∎
We now turn to the proof of Theorem A.3, which we prepare by a series of lemmas.
Lemma A.5.
For every pseudometric $\rho$ we have
  1. (i)
    $E^{\sharp}(\rho)\subseteq\mathbb{R}^{n}_{\geq 0}$;
  2. (ii)
    The complex $E^{\sharp}(\rho)$ is a subcomplex of $E(\rho)$.
Proof.
For the first claim, we show that $E^{\sharp}(\rho)$ has no vertex outside the positive orthant. Consider $x\in E^{\sharp}(\rho)$ and suppose that $x_{1}<0$. Then for every $i,j\neq 1$ we have $x_{i}+x_{j}>x_{i}+x_{1}+x_{1}+x_{j}\geq d_{1i}+d_{1j}\geq d_{ij}$, hence $x_{i}+x_{j}>d_{ij}$. This means that the defining equations of $P^{\sharp}(\rho)$ that hold with equality at $x$ are at most those of the type $x_{1}+x_{i}\geq d_{ij}$. Since the intersection of the hyperplanes $x_{1}+x_{i}=d_{ij}$ has codimension at most $n-2$, the face of $P^{\sharp}(\rho)$ containing $x$ cannot be a vertex.
For the second claim let $F^{\sharp}$ be a face of $E^{\sharp}(\rho)$, i.e., a bounded face of $P^{\sharp}(\rho)$. Surely $F:=F^{\sharp}\cap\mathbb{R}^{n}_{\geq 0}$ is a bounded face of $P(\rho)$, and by claim (i) we have $F=F^{\sharp}$. Thus $F^{\sharp}$ is a bounded face of $P(\rho)$ as well. This show that every face of $E^{\sharp}(\rho)$ is a face of $E(\rho)$. The claim follows. ∎
Lemma A.6.
Let $\rho$ be a pseudometric. Then, for every $k\in[n]$ the following holds.
  • (i)
    For all $k\in[n]$, the point $v^{(k)}$ is a vertex of $E^{\sharp}(\rho)$;
  • (ii)
    The complex $E^{\sharp}(\rho)$ contains the ray $R_{k}:=v^{(k)}+\mathbb{R}_{\geq 0}s^{(k)}$, where
    $s^{(k)}=e_{1}+\ldots e_{k-1}-e_{k}+e_{k+1}+\ldots+e_{n}.$
Proof.
It is straightforward to check that all inequalities that define $P^{\sharp}$ are satisfied by $v^{(k)}$, thus $v^{(k)}\in P^{\sharp}$. We show that $v^{(k)}$ is indeed a vertex, hence a bounded face, of $P^{\sharp}$. Let $i_{0}$ and $j_{0}$ be indices for which the minimum defining $\xi_{k}$ is attained, i.e., $d_{i_{0}k}+d_{kj_{0}}-d_{i_{0}j_{0}}$. Now $v^{(k)}$ lies on the hyperplane $H_{0}$ with equation $x_{i_{0}}+x_{j_{0}}=d_{i_{0}j_{0}}$. Moreover, $v^{(k)}$ is contained in the hyperplanes $H_{ik}:=\{x\in\mathbb{R}^{n}\mid x_{i}+x_{k}=d_{ik}\}$ for all $i\neq k$. The normals of the planes $H_{0}$ and $H_{ik}$ are a linearly independent set of rank $n$ and thus their intersection, since it is nonempty, has dimension 00. ∎
Lemma A.7.
  • (i)
    $E(\rho)\cap H_{k}=\{h^{(k)}_{i}\}$ and in particular every $h^{(k)}$ is a vertex of $E$.
  • (ii)
    If $\xi_{k}>0$ then $E(\rho)\setminus\{v^{(k)}\}$ is disconnected, and the induced partition of the vertices of $E(\rho)$ is a bipartition with $h^{(k)}$ in a singleton block.
Proof.
The first statement is [Dress, Theorem 3.(ii)].
For the second statement notice first that for $i,j\neq k$ we have
$v^{(k)}_{i}+v^{(k)}_{j}=d_{ki}+d_{kj}-2\xi_{k}\geq d_{ki}+d_{kj}-(d_{ki}+d_{kj}-d_{ij})=d_{ij}$.
Moreover, for all $i\neq k$ we have
$v^{(k)}_{i}+v^{(k)}_{k}=d_{ki}-\xi_{k}+\xi_{k}=d_{ki}$
Thus $v^{(k)}$ satisfies the hypotheses of [Dress, Theorem 6] with respect to the bipartition $[n]=\{k\}\uplus([n]\setminus\{k\})$. This theorem says that $E(\rho)\setminus v^{(k)}$ is the union of two open (in $E$) subsets:
$\underbrace{\{x\in E(\rho)\mid x_{k}<v^{(k)}_{k}\}}_{=:A}\uplus\underbrace{\{x\in E(\rho)\mid x_{i}<v^{(k)}_{i}\textrm{ for some }i\in[n]\setminus\{k\}\}}_{=:B}$
If $\xi_{k}>0$ neither subset is empty, since in that case $h^{(k)}\in A$ (because $h^{(k)}_{k}=0<\xi_{k}=v^{(k)}_{k}$) and $h^{(j)}\in B$ for all $j\neq k$ (indeed $v^{(k)}\in E(\rho)$ means $v^{(k)}_{j}=d_{jk}-\xi_{k}\geq 0$ for $j\neq k$. Thus $h^{(j)}_{k}=d_{jk}\geq\xi_{k}=v^{(k)}_{k}$, hence $h^{(j)}\not\in A$, i.e., $h^{(k)}\in B$).
It remains to show that all all-positive vertices of $E(\rho)$ are in $B$. If $n\leq 2$ there is nothing to show. Thus let $n\geq 3$. An all-positive vertex $y$ of $E(\rho)$ is, by Item (ii), a vertex of $E^{\sharp}(\rho)$. Since the intersections of all hyperplanes $x_{i}+x_{k}=d_{ik}$ with $i\neq k$ has codimension at most $n-2$, there is at least a pair $i\neq j$ with $i,j\neq k$ such that $y_{i}+y_{j}=d_{ij}$. Now we can compute
$2y_{k}+d_{ij}=2y_{k}+y_{i}+y_{j}\geq d_{ik}+d_{jk}.$
Thus
$y_{k}\geq\frac{1}{2}(d_{ik}+d_{jk}-d_{ij})\geq\xi_{k}$
hence $y\not\in A$ and so $y\in B$. ∎
Lemma A.8.
If $\xi_{k}>0$, the line segment $e^{(k)}$ joining $h^{(k)}$ $v^{(k)}$ is a cell of $E(\rho)$; moreover, it is the only cell of $E(\rho)$ containing $h^{(k)}$.
Proof.
Notice that, with the notation of Lemma A.6.(ii), we have $h^{(k)}=v^{(k)}+\xi_{k}s^{(k)}$ and, for all $i\neq k$, $v^{(k)}_{i}+\lambda s^{(k)}_{i}>0$ for $0<\lambda<\xi_{k}$. Thus Since $R_{k}$ is a face of $P^{\sharp}(\rho)$, $e^{(k)}=R_{k}\cap\mathbb{R}^{n}_{\geq 0}$ is a face of $P(\rho)$, which is bounded by construction. This proves the first claim.
For the second claim assume by way of contradiction that $h^{(k)}$ is contained in another cell. Then, this cell would have to have at least another vertex $w$ (different from $v^{(k)}$), so even after removing $v^{(k)}$ there would be a path in $E(\rho)$ connecting $h^{(k)}$ with $w$. But this contradicts the disconnectedness claim of Lemma A.7.(ii). ∎
Proof of Theorem A.3.
The first claim is Lemma A.6.(i), the second is Lemma A.5.(ii), the third follows by combining Lemma A.7.(i) and Lemma A.8
Appendix B Table of strict KRW polytopes on $5$ points
In the following table, each row represents one of the $65$ combinatorial types of KRW polytopes of strict metric spaces on $5$ points. For each combinatorial type, we list a sample metric $d$, the $f$-vector of the polytope $KRW(d)$ and the stabilizer subgroup of the cone of $\mathscr{F}_{5}$ containing $d$ (see Section 8.2).
The sample metrics are encoded as vector tuples $(d_{11},d_{12},d_{13},d_{14},d_{15},d_{23},d_{24},d_{25},d_{34},d_{35},d_{45})$. See Table 5 below for the encoding of the stabilizer subgroups.
Table 4. Types of strict metrics for $n=5$.
$f$-vector Nr. sample metric stabilizer
20 60 70 30 1.1 2 2 2 2 2 2 2 2 2 2 $S_{5}$
20 72 94 42 2.1 6 5 5 5 5 5 5 6 6 6 $D_{12}$
2.2 3 3 3 2 2 2 3 2 3 3 $D_{12}$
20 76 102 46 3.1 6 5 5 5 5 5 5 4 6 6 $V_{1}$
20 80 112 52 4.1 8 7 7 5 5 5 7 6 8 8 $C_{2}$
4.2 8 7 6 6 7 6 6 5 7 8 $C_{2}$
4.3 8 7 7 5 5 5 7 4 8 8 $C_{2}$
20 80 114 54 5.1 10 8 7 7 8 7 7 8 8 10 $D_{8}$
5.2 3 4 4 2 3 3 3 2 4 4 $D_{8}$
5.3 4 3 3 3 3 3 3 3 4 4 $V_{1}$
5.4 6 7 7 5 7 7 5 4 8 8 $V_{1}$
20 82 116 54 6.1 6 7 6 5 5 6 5 5 6 7 $C_{2}$
20 84 122 58 7.1 8 7 6 6 5 6 6 5 7 8 $C_{2}$
7.2 8 8 7 6 6 7 6 5 8 9 id
7.3 8 9 7 6 5 7 6 6 7 9 id
7.4 8 9 7 6 7 7 6 6 7 9 id
7.5 8 9 8 5 5 6 7 5 8 9 id
7.6 8 7 9 5 9 7 7 6 8 10 $C_{2}$
7.7 10 9 8 6 7 6 8 5 9 10 id
20 84 124 60 8.1 12 9 8 8 9 8 8 8 10 12 $C_{2}$
8.2 5 4 4 3 3 3 4 3 5 5 $C_{2}$
8.3 8 9 8 6 9 8 6 5 9 10 $C_{2}$
8.4 8 9 9 5 7 7 7 4 10 10 $C_{2}$
20 86 128 62 9.1 12 10 9 7 8 7 9 8 10 12 $C_{2}$
9.2 10 8 7 7 6 7 7 7 9 10 id
9.3 10 9 8 6 5 6 8 5 9 10 $C_{2}$
9.4 10 10 8 7 6 8 7 6 9 11 id
9.5 10 10 8 7 8 8 7 6 9 11 id
9.6 10 11 8 7 7 8 7 7 8 11 id
9.7 10 10 9 6 6 7 8 5 10 11 id
9.8 10 11 9 6 5 7 8 6 9 11 id
9.9 10 11 9 6 7 7 8 6 9 11 id
9.10 30 26 25 17 22 17 25 17 25 30 id
9.11 36 26 25 23 28 23 25 26 28 36 $C_{2}$
9.12 30 23 22 20 25 20 22 17 25 30 id
9.13 8 10 9 5 8 7 7 5 9 10 $C_{2}$
9.14 8 10 9 6 8 9 6 5 10 11 id
9.15 8 11 10 5 7 8 7 5 10 11 $C_{2}$
9.16 10 9 10 6 11 8 8 7 9 12 $C_{2}$
9.17 10 9 11 5 9 7 9 6 10 12 id
9.18 30 23 31 17 31 23 25 20 28 36 id
9.19 10 12 9 6 8 11 8 7 12 11 $C_{2}$
9.20 30 25 20 22 17 22 20 17 25 30 $C_{2}$
20 88 134 66 10.1 14 11 10 8 9 8 10 8 12 14 id
10.2 14 11 9 9 9 9 9 9 11 14 $C_{2}$
10.3 12 10 9 7 6 7 9 7 11 12 id
10.4 12 12 9 8 8 9 8 7 10 13 id
10.5 12 12 10 7 6 8 9 6 11 13 id
10.6 12 12 10 7 8 8 9 6 11 13 id
10.7 12 13 10 7 7 8 9 7 10 13 id
10.8 42 32 31 23 28 23 31 26 34 42 id
10.9 36 29 28 20 25 20 28 17 31 36 id
10.10 42 29 28 26 31 26 28 26 34 42 id
10.11 10 11 10 6 9 8 8 5 11 12 id
10.12 10 12 10 7 10 10 7 6 11 13 id
10.13 10 12 11 6 8 9 8 5 12 13 id
10.14 10 13 11 6 9 9 8 6 11 13 id
10.15 12 11 12 6 11 8 10 7 11 14 id
10.16 10 11 13 5 11 9 9 6 12 14 $C_{2}$
10.17 36 29 34 20 37 26 28 23 31 42 id
10.18 36 29 37 17 31 23 31 20 34 42 id
10.19 42 46 32 23 26 40 31 26 45 43 $C_{2}$
10.20 12 14 11 6 8 11 10 7 14 13 id
10.21 36 28 23 25 20 25 23 23 31 36 id
10.22 42 46 32 31 26 40 23 26 37 51 $C_{2}$
[…]
trivial 1 ()()( ) id
gen. by double transposition 2 $\langle(15)(24)\rangle$ $C_{2}$
gen. by pair of disjoint transpositions 4 $\langle(12),(34)\rangle$ $V_{1}$
gen. by double transpositions
      on 4 elements 4 $\langle(12)(45),(14)(25)\rangle$ $V_{2}$
dihedral of order $8$ 8 $\langle(1234),(13)\rangle$ $D_{8}$
dihedral of order $10$ 10 $\langle(12345),(25)(34)\rangle$ $D_{10}$
direct product of $S_{3}$ and $S_{2}$ 12 $\langle(123),(12),(45)\rangle$ $D_{12}$
Table 5. List of types of subgroups of $S_{5}$ appearing as stabilizer of chambers in $\mathcal{W}_{5}$.